### v1

# Soft-CAM: Making black box models self-explainable for high-stakes decisions

Abstract
Convolutional neural networks (CNNs) are widely used for high-stakes applications like medicine, often surpassing human performance. However, most explanation methods rely on post-hoc attribution, approximating the decision-making process of already trained black-box models. These methods are often sensitive, unreliable, and fail to reflect true model reasoning, limiting their trustworthiness in critical applications. In this work, we introduce SoftCAM, a straightforward yet effective approach that makes standard CNN architectures inherently interpretable. By removing the global average pooling layer and replacing the fully connected classification layer with a convolution-based class evidence layer, SoftCAM preserves spatial information and produces explicit class activation maps that form the basis of the model’s predictions. Evaluated on three medical datasets, SoftCAM maintains classification performance while significantly improving both the qualitative and quantitative explanation compared to existing post-hoc methods. Our results demonstrate that CNNs can be inherently interpretable without compromising performance, advancing the development of self-explainable deep learning for high-stakes decision-making.
1 Introduction
Convolutional Neural Networks (CNNs) have revolutionized computer vision by efficiently capturing local patterns, reducing parameters, and accelerating convergence, enabling superior performance in tasks like image recognition and object detection [35, 40]. However, their lack of interpretability limits adoption in high-stakes fields like medicine, where transparency and trust are crucial. To explain CNNs, numerous saliency-based methods have been proposed, including class activation maps (CAM) [61] and their variants [14, 47, 56, 59], gradient-based techniques [48, 50, 51, 53], and even perturbation- or occlusion-based methods [22, 56, 59]. These methods have been widely adopted to explain the decisions of trained black-box models.
Such saliency map-based techniques offer explanations for CNN classifiers that claim to highlight regions in the input image most relevant to the model’s prediction. These explanations are generated post-hoc, typically after a model is trained [10, 27]. Studies have shown significant limitations in their effectiveness, especially in clinical settings [5]. Post-hoc saliency methods often lack faithfulness, reliability, and consistency, resulting in explanations that may not accurately reflect the model’s decision-making process [3, 58]. Moreover, they struggle to localize relevant regions in medical imaging [4], where the limited availability of ground-truth annotations makes it difficult to assess their trustworthiness. To overcome these challenges, inherently and/or self-explainable models have been introduced [45], designed explicitly to provide interpretable insights by incorporating explanations within their architecture [11, 13, 16, 18, 52]. These models generate more trustworthy and faithful explanations that align closely with the model’s actual reasoning [28]. However, self-explainable models generally use specific architectures [13, 16, 52], which limits their applicability and generalization to widely used CNN architectures.
Motivated by these challenges, we propose SoftCAM, a straightforward generalization of Class Activation Maps (CAM) that uses a convolution-based classifier to transform any black-box CNN into a self-explainable model. By removing the final pooling layer and replacing the fully connected classification layer with 1x1 convolutions, SoftCAM turns classical CNNs into fully convolutional networks, generating class-specific evidence maps that are directly used for predictions. Our contributions are:
  • We introduced SoftCAM, a simple modification to CNNs that enables self-explainability, and experimentally demonstrate that the resulting model preserves classification performance across three clinically relevant medical datasets spanning different imaging modalities.
  • We showed that regularizing evidence maps using ElasticNet, a regularizer combining both ridge and lasso penalties, enhances the model’s explanations.
  • We evaluated five widely used traditional CAM-based post-hoc explanation methods, showing that SoftCAM most often outperforms them across a broad range of explainability metrics and across all three considered medical imaging datasets and modalities.
2 Related work
Deep neural networks (DNNs) are widely used in a variety of fields. However, regulatory frameworks such as the European AI Act require AI-based decisions to be explainable to ensure fairness, transparency, and accountability, allowing users to understand their decision-making process [2, 41]. Explainability can help verify and even improve performance by detecting shortcuts and identifying clinically relevant features, ultimately fostering greater trust in decision-making, especially in fields like healthcare [17].
Existing explainable AI (XAI) methods for image analysis can be broadly categorized into attribution-based and non-attribution-based approaches [10, 27, 57]. Attribution-based methods explain “where” important features exist in the input by generating saliency or heatmaps that assign importance scores to individual pixels or regions, helping visualize their contribution to the model’s decision. These include perturbation-based methods [29] and class activation maps [26], which consist of both gradient-based [47, 51, 53] and gradient-free approaches [44, 56, 61]. In contrast, non-attribution-based approaches explain “why” a decision was made without relying on importance scores, instead using techniques such as concept-based (ACE [23], TCAV [32], CBM [33]), prototype-based [15, 16], or counterfactual-based [7, 12, 24, 52] methods to analyze model behavior from different perspectives. These approaches also differ in how explanations are obtained and which architectures they can be applied to. Attribution-based methods typically provide post-hoc, local explanations by offering input-specific insights into black-box CNN models after training. In contrast, non-attribution-based methods are generally inherently interpretable by design, promoting transparency and enabling global understanding of the model’s decision-making process across the entire dataset [45]. However, in some cases, self-explainable models may be less effective for complex tasks, highlighting a tradeoff between interpretability and performance, where increasing transparency may sometimes come at the cost of classification performance [57].
Our method, SoftCAM, relies on an explicit class-evidence layer based on convolutional operations for classification, offering class-specific, attribution-based explanations while remaining inherently interpretable, unlike post-hoc approaches. Moreover, it maintains predictive performance comparable to its corresponding non-interpretable black-box models. Closely related work includes [6] and [18]. In [6], a dual-branch approach is used, where one branch leverages a traditional black-box CNN model (ResNet-18) for classification, and the second branch uses weight-sharing for post-hoc explanations, requiring two forward passes for inference. In the second branch for post-hoc explanation, the global average pooling layer (GAP) is removed, and the linear classifier is replaced by convolutional layers that share weights during inference to generate class-specific activation maps. In contrast, SoftCAM is trained end-to-end, providing both predictions and explanations in a single forward pass, eliminating the need for additional computational overhead or weight sharing. Furthermore, while [18] uses explicit class-evidence maps to enhance the explanation of a self-explainable bag-of-local-feature model (BagNet [13]), our method transforms black-box models into self-explainable models. We evaluated our approach on a range of medical datasets, comparing the resulting explanations to various post-hoc attribution-based methods, including both gradient-free and gradient-based techniques.
3 Method
Preliminaries
Given an input image $\mathbf{X}\in\mathbb{R}^{H_{X}\times W_{X}\times C_{X}}$ with height $H_{X}$, width $W_{X}$, and the number of channels $C_{X}$, consider a CNN network $f_{\theta}$ that maps $\mathbf{X}$ to a probability distribution $\mathbf{\hat{y}}=f_{\theta}(\mathbf{X})\in\mathbb{R}^{C}$, where $C$ is the number of classes, and $y^{c}\in\mathbf{y}$ represents the predicted probability for class $c$. The network consists of a feature extractor $g_{\phi}$, and a classifier layer $h_{\psi}$, with learnable parameters $\phi$ and $\psi$. The feature extractor generates a feature map $\mathbf{Z}=g_{\phi}(\mathbf{X})\in\mathbb{R}^{N\times M\times D}$, where $N\times M$ denotes the spatial size and $D$ is the feature dimension (e.g., $D=2048$ for most ResNet variants). The classifier then predicts the final output based on the extracted features. Let $\mathcal{A}=\{\mathbf{A}_{k}\}_{k=1}^{D}$, the set of activation maps obtained from the feature extractor, where $A_{k}$ is the activation of the $k$-th neuron. Let, $S^{c}_{\text{Map}}\in\mathbb{R}^{N\times M}$ be the 2D saliency map, providing a visual explanation of the model’s prediction for class $c$. This paper explores how to train self-explainable CNNs to simultaneously generate both the prediction $y^{c}$ and its corresponding explanation $S^{c}_{\text{Map}}$.
Traditional CNN architectures employ a GAP layer to reduce the feature map to $1\times D$, followed by a classification module consisting of one or more linear fully connected layers (FCLs) to generate the final prediction. Post-hoc methods are then typically used to explain the model’s decision.
3.1 CAM-based methods
Class Activation Maps (CAMs) [61] are closely related to our approach, offering local visual explanations of CNN predictions by generating saliency maps for individual inputs. CAM achieves this by linearly combining the feature maps from the final convolutional layers with importance coefficients from the FCL classifier, thereby producing class-wise attribution maps as follows:
$S_{\text{CAM}}^{c}(x_{1},x_{2})=\sum_{k=1}^{D}w_{k}^{c}A_{k}(x_{1},x_{2}),$ (1)
where $A_{k}(x_{1},x_{2})$ is the activation of neuron $k$ in the feature map at spatial location $(x_{1},x_{2})$, and $w_{k}^{c}$ denotes the importance weight associated with class $c$ for unit $k$ in the fully connected layer.
Originally, CAM was designed for CNNs with GAP and FCL, but has been extended to gradient-based methods using class score gradients to compute importance weights [14, 47, 48]. This extension enabled CAM-based techniques to be applied to a broader range of CNN architectures, particularly those where the GAP layer is followed by multiple FCLs, as seen in models like VGG [49] and InceptionV3 [54]. For example, GradCAM [47] extends the original CAM approach by backpropagating the gradient from a target class to the input layer to highlight the image regions that strongly influence the model’s prediction. GradCAM is formulated as
$S^{c}_{\text{Grad-CAM}}(x_{1},x_{2})=\text{ReLU}\Bigg{(}\sum_{k=1}^{D}w_{k}^{c}A_{k}(x_{1},x_{2})\Bigg{)},$ (2)
where the weight coefficients are computed as $w_{k}^{c}=\frac{1}{N\times M}\sum_{i}^{N}\sum_{j}^{N}\frac{\partial y^{c}}{\partial A_{k}(i,j)}$. Here, $A_{k}(i,j)$ is the activation value at location $(i,j)$ on $A_{k}$, and the rectified linear unit (ReLU) is applied to ensure that the final activation map considers only the features that positively influence class $c$. Following GradCAM, several variations have been proposed, including gradient-based approaches such as SmoothGrad [50], GradCAM++ [14], guided-backpropagation [51], and integrated gradients [53], as well as gradient-free methods like ScoreCAM [56], LayerCAM [30], and OptiCAM [59]. Gradient-based methods primarily differ in how gradients are aggregated to compute importance weights, while gradient-free methods mainly vary in how the weights are computed.
Despite the success of class activation map-based methods in explaining CNN classifiers, including medical applications [8], they have a key limitation: they rely on already trained models and provide post-hoc explanations, which may not accurately reflect the model’s true decision-making process. Additionally, gradient-based methods face inherent challenges such as gradient saturation, where DNN gradients tend to diminish, and false confidence, where the highest activation map weight does not necessarily correspond to the greatest increase in confidence [56]. On the other hand, gradient-free methods are computationally and memory-intensive, as they often require multiple forward passes on perturbed inputs. Finally, CAM-based methods are easy to implement in CNNs with clearly defined spatial feature maps, but face challenges in multi-branch architectures like InceptionV3 due to the complexity of integrating diverse feature maps from parallel convolutional paths.
Figure 1: Overview of softCAM for making black-box CNNs inherently interpretable. (a) Input image. (b) The CNN backbone consists of all layers before the global average pooling layer. (c) Feature map generated by the backbone. (d) Classifier module with $C$ convolutional kernels of size $1\times 1$. (e) Self-explainable class activation maps $\mathbf{A}$, obtained from the classifier with ElasticNet penalty applied to it to enhance interpretability. (f) Final predictions are derived directly from the evidence maps via spatial average pooling followed by the softmax function. Class-specific evidence maps (g) are upsampled and overlaid on the input to visualize the model’s decision-making process.
3.2 Improving CAMs for self-explanability
Motivated by the limitations of post-hoc class activation map-based methods in interpreting CNN classifiers, we introduce SoftCAM (Fig. 1), a straightforward modification of black-box CNN classifiers that makes them self-explainable and inherently interpretable. SoftCAM achieves this by replacing the fully connected classification layer in classical CNNs with an explicit class-evidence convolutional layer, preserving spatial information and providing explanations in a single forward pass, eliminating the need and computational overhead for post-hoc techniques.
We make black-box CNN architectures self-explainable by modifying how predictions are obtained. Any FCL of size $b_{1}\times b_{2}$, where $b_{1}$ and $b_{2}$ denote the number of input and output features, respectively, can be equivalently expressed as a $1\times 1$ convolutional layer with $b_{1}$ input channels and $b_{2}$ output channels [18]. This allows us to replace FCL classifiers in standard CNN architectures with convolutions, removing the GAP layer before classification, while preserving model complexity and spatial localization. The new classifier module $h$ consists of convolutional layers (Fig. 1d) with $C$ convolution kernels of size $1\times 1$ and unit stride, producing class evidence maps (Fig. 1e)
$\mathbf{A}=h_{\psi}(\mathbf{Z})\in\mathbb{R}^{M\times N\times C},$ (3)
where $\psi$ is a learnable parameter. Indeed, $h_{\psi}$ can be viewed as an explainable, soft generalization of classical post-hoc attribution methods (Eq. 1, 2), mapping the low-dimensional feature volume $\mathbf{Z}$ into an interpretable, class-wise activation volume $\mathbf{A}$ whose reduced channel dimension corresponds to the number of target classes. Unlike CAM (Eq. 1) and GradCAM (Eq. 2), which generate post-hoc heuristic explanations, our approach leverages the feature map volume from the backbone and applies a parameterized function $h_{\psi}$ to directly produce class activation maps that are used for prediction. In contrast to classical CAM-based methods, the importance weights are not explicitly defined but are implicitly learned and encoded within the classifier’s parameters.
The resulting architecture is a fully convolutional, self-explainable model, where the final predicted probabilities are computed from the evidence map (Fig.1e), without introducing additional parameters:
$\mathbf{\hat{y}}=\text{Softmax}\bigg{(}\mbox{AvgPool}\Big{(}h_{\psi}\big{(}g_{\Phi}(\mathbf{X})\big{)}\Big{)}\bigg{)}\in\mathbb{R}^{1\times C}.$ (4)
Additionally, the class evidence maps $\mathbf{A}$ serve as built-in explanations, directly representing the contribution of individual input regions to the final prediction (Fig. 1g). Replacing linear FCL layers with convolutional operations offers several advantages. Due to the shift-invariance and position-agnostic properties of CNNs, all image regions are weighted equally when forming the final classification (Fig. 1f). As a result, input feature patches with high activations in the evidence maps contribute most significantly and linearly to the prediction. This behavior mirrors that of simple linear models, where each value in the activation map has a direct and interpretable impact on the output.
3.3 Regularizing SoftCAM for interpretability
By using explicit class-evidence maps, the model can be trained directly with regularization applied to these maps to enhance interpretability. In practice, we apply an ElasticNet regularization [62], which linearly combines the $\ell_{1}$ (Lasso) and $\ell_{2}$ (Ridge) penalties, leading to the following loss function:
$\mathcal{L}(\mathbf{y},\mathbf{\hat{y}})=\mbox{CE}(\mathbf{y},\mathbf{\hat{y}})+\lambda_{1}\sum_{i,j,c}|\mathbf{A}_{c}^{ij}|+\lambda_{2}\sum_{i,j,c}||\mathbf{A}_{c}^{ij}||_{2}.$ (5)
Here, CE denotes the cross-entropy loss, and $\mathbf{y}$ represents the reference class labels. When $\lambda_{2}=0$, the ElasticNet penalty becomes the Lasso penalty, which promotes sparsity in the class evidence maps [18] by removing less informative activations, making it particularly useful for tasks where precision in explanations is crucial. In contrast, when $\lambda_{1}=0$, the ElasticNet reduces to the Ridge penalty, which reduces irrelevant activations without forcing them to zero, which is beneficial when minimizing false negatives is a priority. ElasticNet thus provides a balance between Lasso and Ridge penalties, balancing sparsity and smoothness in the resulting activation maps.
Visualizing explanations.
The evidence map generated by SoftCAM is upsampled to the input resolution for visualization (Fig. 1g). Like most CAM-based methods, such as GradCAM [47], ScoreCAM [56], and LayerCAM [30] that operate on the final convolutional layer, SoftCAM’s explanations are limited by the resolution of the backbone (e.g., 16×16 for VGG-16/ResNet-50 with 512×512 input) due to pooling and striding, leading to lower-resolution saliency maps. However, by introducing the class evidence and classification layer directly atop features, SoftCAM regularizes the evidence map, making it less coarse and thereby enhancing localization. In contrast, gradient-based methods like Integrated Gradients [53] and Guided Backpropagation [51] produce high-resolution saliency maps by computing pixel-level gradients, which may lead to noisy maps, especially when the region of interest spans a broader area, as commonly observed in Chest X-ray images.
Comparison with other approaches.
Unlike post-hoc attribution-based approaches, our method is inherently interpretable from the classification layer and maintains performance comparable to its black-box counterpart, without a significant trade-off, even when regularization is applied to enhance explainability. Compared to [18], our method extends from the concept of interpretable bag-of-local models to general black-box CNN architectures and generalizes the regularization from Lasso to ElasticNet, with extensive evaluations across multiple datasets using a broad range of explanability metrics. Compared to [6], our method is trained end-to-end and does not require post-hoc processing, weight sharing between branches, or an additional forward pass to generate explanations.
4 Experimental setup
Datasets.
We evaluated our approach on three publicly available medical datasets spanning three imaging modalities: the Kaggle Diabetic Retinopathy (DR) [20], Retinal OCT [31], and the RSNA Chest X-Ray (CXR) [1]. The first dataset comprised high-resolution retinal color fundus images, each labeled with a DR severity score ranging from 00 (No DR) to $4$ (Proliferative DR). The second dataset included retinal OCT B-scans images categorized into Drusen, Diabetic macular edema, Choroidal neovascularization, and Normal cases. The final dataset consisted of high-resolution frontal-view chest radiographs labeled for pneumonia detection, with bounding box annotations for pneumonia cases. Additionally, lesion annotations were obtained for $65$ DR images from the Kaggle dataset [17] and $40$ Drusen images from the retinal OCT dataset [16]. Each dataset was split into training, validation, and test sets using different dataset-specific train-validation-test proportions, ensuring that all samples from the same patient were assigned to the same split to prevent data leakage. For full details, see Appendix A.1.
Baseline models.
The effectiveness of our method was evaluated using two widely used black-box CNN architectures: ResNet-50 [25] and VGG-16 [49]. They differ primarily in the design of their classification heads, where ResNet employs a single fully connected layer, while VGG uses multiple. In both models, we explicitly replaced the classification head with our convolutional evidence map layer, adapting the architecture to enable interpretability (see Appendix A.2). The models were sourced from Torchvision [36], initialized with pre-trained weights from ImageNet, and fine-tuned using a consistent setup [^1] Our code with datasets is available at https://anonymous.4open.science/r/SoftCAM-E1A3/. Training was performed over $70$ epochs with a mini-batch size of $16$ on an NVIDIA A40 GPU using PyTorch [42]. A range of data augmentation and preprocessing techniques was applied (see Appendix A.3). For complete training details, see Appendix A.4.
Baseline CAM-based methods.
We qualitatively and quantitatively assessed the explanations generated by our method (SoftCAM) against post-hoc explanation techniques from several state-of-the-art class attribution map-based methods, applied to their respective black-box models. Specifically, we compared our approach with gradient-based methods, including GradCAM [47], Integrated Gradient (Itgd Grad.) [53], Guided Backpropagation (Guided BP) [51], as well as gradient-free methods such as ScoreCAM [56] and LayerCAM ([30]). Guided BP and Itgd Grad. have consistently performed well in producing saliency maps for explaining black-box CNN classifiers on retinal images [8, 17], while GradCAM has shown strong localization performance for chest X-ray interpretation [46]. Gradient-based methods were implemented from Captum [34], whereas gradient-free methods were implemented via TorchCAM [21]. For full descriptions of these methods, see Appendix A.5.
Evaluation metrics.
Models were evaluated on both classification performance and explainability. For binary tasks, performance was measured using accuracy and AUC, while for multi-class tasks, accuracy and the quadratic Cohen’s kappa score were used. AUC reflects class separability, whereas the kappa score captures agreement beyond chance. To assess explainability, we employed several quantitative metrics: Top-k localization precision [18], activation precision [9, 43], activation consistency [18], and faithfulness [16, 39]. We further extended activation precision to define activation sensitivity. For full descriptions of the explainability metrics, see Appendix B.
5 Results
5.1 Making black box CNNs self-explainable maintains classification performance
Table 1: Classification performance for binary disease detection on the test sets. We denote the SoftCAM versions of ResNet and VGG with a .
Kaggle Fundus OCT retinal RSNA CXR
Binary Multi-class Binary Multi-class Binary
Acc. AUC Acc. $\kappa$ Acc. AUC Acc. $\kappa$ Acc. AUC
VGG-16 $0.907$ $0.938$ $0.863$ $0.835$ $0.994$ $1.0$ $0.967$ $0.955$ $0.952$ $0.989$
dense VGG $0.915$ $0.942$ $0.861$ $0.834$ $0.994$ $1.0$ $0.963$ $0.947$ $0.957$ $0.999$
sparse VGG $0.911$ $0.938$ $0.859$ $0.827$ $0.988$ $0.999$ $0.947$ $0.929$ $0.953$ $0.990$
ResNet-50 $0.899$ $0.923$ $0.850$ $0.800$ $0.994$ $0.999$ $0.970$ $0.963$ $0.953$ $0.988$
dense ResNet $0.899$ $0.926$ $0.851$ $0.811$ $0.994$ $1.0$ $0.974$ $0.960$ $0.942$ $0.983$
sparse ResNet $0.895$ $0.923$ $0.851$ $0.801$ $0.996$ $1.0$ $0.963$ $0.955$ $0.941$ $0.979$
We first evaluated our method on clinically relevant classification tasks, including retinal disease classification from color fundus and OCT retinal images, as well as pneumonia detection from chest X-rays. For the fundus and OCT retinal datasets, both binary classification ({0} vs. {1-4} for fundus and Normal vs. Drusen for OCT) and multi-class classification tasks were considered, as reference labels were available. In contrast, the RSNA CXR dataset only included labels for pneumonia detection, restricting the evaluation to the binary task. For each CNN architecture, the “dense” model corresponds to our method without regularization ($\lambda_{1}=\lambda_{2}=0$), while the “sparse” model is obtained by applying a lasso penalty ($\lambda_{2}=0$) and choosing an appropriate value for $\lambda_{1}$ (e.g. $\lambda_{1}=1.10^{-6}$ for VGG and $\lambda_{1}=5.10^{-5}$ for ResNet on the fundus dataset). The sparsity parameter was selected by balancing classification accuracy and AUC on the validation set (see Appendix C).
Our results show that SoftCAM models, which use explicit self-explainable class evidence maps, preserve classification performance comparable to their corresponding black-box counterparts (Tab. 1). Moreover, introducing the Lasso regularization penalty on the class evidence map did not significantly degrade performance; in some cases, it even led to slight improvement. These findings suggest that using convolutional layers in the classification head is an effective and promising approach for developing high-performing, self-explainable CNN models.
5.2 SoftCAM provides inherently interpretable visual explanations
Figure 2: Example explanations generated by different methods from ResNet-50. The first column shows disease images with reference annotations, indicated by green markers or bounding boxes. Each row, from top to bottom, corresponds to fundus, OCT, and Chest X-ray images, respectively. The next five columns present saliency maps generated by post-hoc explanation methods, gradient-free (ScoreCAM, LayerCAM) and gradient-based (GradCAM, Guided BP, Itgd Grad). The final two columns showcase our proposed inherently interpretable dense and sparse SoftCAM explanations.
We qualitatively compared the evidence maps of SoftCAM variants with saliency maps generated by the five state-of-the-art CAM-based methods. Overall, our method produced more visually interpretable maps with high evidence regions centered on annotated lesions (Fig. 2). We observed that the regions highlighted by the sparse SoftCAM models are mostly a subset of those identified by the dense SoftCAM, reflecting the effect of the sparsity constraint. Additional results, including those for VGG-16 and other illustrative examples, are provided in Appendix D.1.
On healthy images, sparse SoftCAM evidence maps exhibited overall more negative activations, in contrast to the positive activations observed on disease images. To assess this quantitatively, we computed the activation consistency [18], calculating the proportion of positive and negative activations across disease and healthy samples. These findings were consistent with the qualitative observations (e.g. dense vs. sparse SoftCAM on the fundus dataset using ResNet: $0.55$ vs. $0.27$ for the proportion of positive activation on disease images). For full analysis, see Appendix D.2.
5.3 SoftCAM provides localized and faithful explanations
Figure 3: Quantitative evaluation of explanations generated by different methods. The first row shows the localization precision of the saliency maps on the Fundus and OCT datasets, evaluated against their respective ground truth. The second row presents the sensitivity analysis assessing the faithfulness of the generated explanations. Columns a,b show ResNet results, while c,d correspond to VGG. Higher precision means better localization; lower sensitivity implies more reliable explanations.
[…]
5.4 Ridge regularization improves explanation for large disease regions
Since the CXR dataset provided larger bounding boxes localizing disease regions, unlike the point-wise lesion annotations available in the fundus and OCT datasets, we computed activation precision [9, 43], which measures the proportion of the class-guided explanation that fall within the ground-truth bounding boxes, emphasizing precision by penalizing only false positives. However, it does not account for sensitivity or penalize false negatives. To address this limitation, we extended this metric to activation sensitivity (see Appendix B.2), which penalizes false negatives to better assess the explanation completeness, especially important in clinical imaging tasks where missing relevant regions can be critical, such as in multi-focal infectious diseases like pneumonia like pneumonia [37]. We further investigated how different regularization strategies affect explanation quality. While Lasso regularization promoted sparsity by shrinking some activations to zeros, ridge regularization encouraged small (but nonzero) values, resulting in denser evidence maps. To evaluate this, we trained a ridge SoftCAM model ($\lambda_{1}=0$) and compared its performance to dense and sparse SoftCAM, as well as to the post-hoc explanation methods. The ridge penalty values were selected to balance classification performance ($\lambda_{2}=7.10^{-5}$ vs. $\lambda_{2}=2.10^{-4}$ for ResNet and VGG; see Appendix E.1).
Under comparable classification performance (Acc.=$0.95$ for Ridge ResNet*, and VGG*), we found that all SoftCAM variants (dense, sparse, and ridge) generally outperformed the evaluated posthoc methods in both activation precision and activation sensitivity (Fig. 4; Appendix E.2, E.3). Specifically, sparse SoftCAM achieved the highest activation precision, while ridge SoftCAM excelled in activation sensitivity. Dense SoftCAM consistently performed in between, underscoring the importance of balancing lasso and ridge regularization via ElasticNet to adapt to varying interpretability needs.
Figure 4: Example of localization evaluation on the CXR dataset for pneumonia detection. The first row shows saliency maps generated by different methods from the ResNet model, and the second row from the VGG model. Ground-truth bounding boxes are overlaid on each map, with the top-right value indicating the activation precision, while the top-left value indicates the activation sensitivity.
5.5 SoftCAM provides resource-efficient and faithful explanations for multi-class tasks
Finally, we extended our method to the multi-class setting for retinal disease diagnosis. We retrained the same training setup as for the binary tasks, adjusting the output classes in the evidence layer to $5$ for DR grading (fundus dataset) and 4 for retinal disease classification (OCT dataset). Given the small size of retinal lesions, we used Lasso regularization, selecting $\lambda_{1}$ values that balanced performance (e.g. $\lambda_{1}=9.10^{-4}$ vs. $\lambda_{1}=3.10^{-6}$ for ResNet and VGG on the OCT dataset; Appendix F.1). Both dense and sparse models achieved performance comparable to their respective black-box baselines (Tab.1), with a slight improvement in Kappa on the fundus dataset when using the ResNet backbone.
As no ground-truth lesion annotations were available for the multi-class tasks, we evaluated the faithfulness of the explanations by measuring their contribution to model predictions. For correctly classified test samples, we progressively removed top-k ($k=30$) ranked patches (based on the explanation maps; see Sec.5.3) and tracked the average drop in class confidence. In both tasks, the dense and sparse SoftCAM achieved superior performance, with sparse SoftCAM yielding the lowest area under the deletion curve, indicating the highest faithfulness (see Appendix F.2, F.3).
Notably, the sparse SoftCAM produced class-wise explanations that aligned well with class model confidence, showing minimal evidence in healthy classes (Fig.5; Appendix F.4, F.5 for VGG and more examples). In the case of DR detection, a progressive disease, it is expected that images labeled with grade $x$, where $51<x<$ 5, may still exhibit features from earlier stages, consistent with explanations. Unlike post-hoc CAM-based methods, which require backpropagation or perturbation for each class, SoftCAM generates class-specific explanations during prediction in a single forward pass, making it more resource-efficient.
Figure 5: Examples of multi-class explanations using ResNet. For a severe DR example from the Kaggle dataset, the first row shows class-specific dense SoftCAM evidence map explanations, while the second presents explanations from the sparse SoftCAM.
6 Discussion
Here, we introduced SoftCAM, a straightforward yet effective approach for transforming black-box CNN models into inherently interpretable architectures. We tested SoftCAM on a diverse range of medical imaging tasks, including color fundus photographs, retinal OCT scans, and Chest X-rays for disease diagnosis. Importantly, SoftCAM-variants maintained performance comparable to that of the original CNN models for classification and generally outperformed post-hoc explainability techniques. SoftCAM produces explicit class evidence maps that directly contribute to the model’s prediction. This integration enables single forward-pass generation of explanations aligned with the classification output, resulting in resource-efficient and self-explainable CNNs.
We evaluated our method for two widely used CNN backbones: ResNet-50 and VGG-16, assessing both classification performance and explainability. Despite some differences, both ResNet and VGG models employ large receptive fields, resulting in low-resolution feature maps. Consequently, the class-evidence layer operates on coarse feature maps, producing coarse-grained explanations. In the future, we could explore the integration of SoftCAM with other standard architectures like ViT [19].
Our work presents a major step forward in the development of powerful self-explainable models, demonstrating that interpretable-by-design architectures can preserve, and in some cases even improve upon, the classification performance of state-of-the-art models by modifying standard, well-tested CNN architectures without the need for complicated additional concepts such as prototypes [15]. Beyond performance, SoftCAM provides deeper insights into the model-decision-making process, offering a powerful tool for understanding mistakes and detecting spurious correlations, without relying on widely used post-hoc explanation methods. By leveraging ElasticNet regularization, which is task-specific, user can flexibly balance localization precision and sensitivity according to their application needs. This is especially relevant for CNN-based classifiers deployed in high-stakes decision-making contexts. We hope this contribution will pave the way toward designing more accurate and interpretable CNN models, ultimately fostering trust, adoption, and integration in critical real-world settings such as in medicine.
Acknowledgments and Disclosure of Funding
This project was supported by the Hertie Foundation, the German Science Foundation (Excellence Cluster EXC 2064 “Machine Learning—New Perspectives for Science”, project number 390727645). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting KD.
Appendix A Implementation Details
A.1 Datasets
We evaluate our approach on three publicly available medical imaging datasets spanning three different modalities: the Kaggle Diabetic Retinopathy (DR) [20], the Retinal OCT dataset [31], and the RSNA Chest X-ray (CXR) dataset [1].
  • Kaggle DR Dataset. This dataset comprises $88,702$ high-resolution retinal fundus images labeled for DR severity on a 5-point scale from 00 (No DR) to $4$ (Proliferative DR). After applying an automated quality filtering pipeline using an ensemble of EfficientNet models [55] trained on the ISBI2020222https://isbi.deepdr.org/challenge2.html challenge dataset, we retained $45,923$ images from $28,984$ subjects. The resulting class distribution was $73\%$ (class 0), $15\%$, $8\%$, $3\%$, and $1\%$. For binary classification (early DR detection), we grouped class {0} vs. {1,2,3,4}, yielding an imbalance of $73\%$ vs. $27\%$. Additionally, lesion annotations for $65$ images were obtained from [17] for evaluating the model’s explanations at localizing DR-related lesions.
  • Retinal OCT Dataset. This dataset consists of $108,315$ B-scans categorized into four classes: Drusen, Diabetic macular edema (DME), Choroidal neovascularization (CNV), and Normal. A separate test set of $1,000$ B-scans is provided. Following [16], we excluded low-resolution scans (width $\leq$ 496). As preliminary experiments showed that using the full dataset did not significantly improve performance, we subsampled the training set (by randomly removing half of the healthy images [16]) to $34,962$ scans ($8,616$ Drusen, $26,346$ Normal) for binary classification (Drusen vs. Normal), preserving the original class imbalance ($73\%$ vs $27\%$). Additionally, we used $40$ drusen-annotated B-scans from [16] to evaluate the model’s explanations at localizing drusen lesions. For the multi-class classification task, the training was randomly reduced to $17,200$ images while maintaining the original class distribution: $45\%$ Normal, $34\%$ CNV, $10\%$ DME, and $9\%$ Drusen.
  • RSNA Chest X-ray Dataset. This dataset includes $30,227$ frontal-view chest radiographs labeled as “Normal”, “No Opacity/Not Normal”, and “Opacity” (indicative of pneumonia). Pneumonia cases come with bounding box annotations, which facilitate the evaluation of the model’s explanations. For our binary classification task, we selected images labeled as either “Normal” or “Opacity”, resulting in $14,863$ images with a $60\%$ vs. $40\%$ class distribution.
Each dataset was split into training ($75\%$), validation ($10\%$), and test ($15\%$) sets, except for the Retinal OCT dataset, which followed an $80\%\text{-}20\%$ training-validation split, due to its predefined test set ( $250$ images per class). All training splits used in our experiments are provided in CSV format and publicly available via the project’s GitHub [^3] Code and CSV files are available at https://anonymous.4open.science/r/SoftCAM-E1A3/ repository.
A.2 Baseline models
The effectiveness of our method was evaluated using two widely adopted black-box CNN architectures: ResNet-50 [25] and VGG-16 [49]. These models were chosen due to their distinct architectures, such as depth, theoretical receptive field size, and classification head design, which allow for a broad assessment of our method’s generalizability. In both models, the standard classification head was replaced with our proposed convolutional evidence map layer to enable inherent interpretability. For ResNet50, we removed the global average pooling layer and final fully connected layer, substituting them with a class evidence layer consisting of $C$ convolutional filters ($1\times 1$, stride 1), where $C$ is the number of output classes. This layer directly produces class-specific evidence maps (Sec. 3.2).
For VGG-16, which uses a series of fully connected layers (FCLs) in its classifier head, each FCL was replaced by an equivalent $1\times 1$ convolutional layer. Specifically, an FCL of size $b_{1}\times b_{2}$ was transformed into a convolutional layer of size $b_{1}\times b_{2}\times 1\ times1$, preserving the original parameter count and model capacity. These architectural adjustments maintain model complexity and capacity while introducing interpretability directly into the classification mechanism.
A.3 Data preprocessing
Fundus images were preprocessed by cropping them to a square shape using a circle-fitting method as described in [38]. All datasets were then resized to $512\times 512$ pixels, except for the retinal OCT dataset, which was resized to $496\times 496$ to better match its original lower resolution. Image intensities were normalized using the mean and standard deviation computed from the respective training sets.
During training, consistent data augmentation strategies were applied across all datasets. These included flipping, rotation, random cropping, and translation, each applied with a fixed probability. For the Kaggle dataset, which contains color fundus images, additional color augmentations were introduced to improve generalization.
A.4 Training setup
All models were obtained from Torchvision and initialized with pretrained ImageNet weights. They were subsequently fine-tuned on each dataset using a consistent training setup. Following [16, 18], we employed the cross-entropy loss function and optimized model parameters using stochastic gradient descent (SGD) with Nesterov momentum (momentum factor of $0.9$). The initial learning rate was set to $1.10\text{-}3$, and a clipped cosine annealing learning rate scheduling was applied with the minimum learning rate set to $1.10\text{-}4$. Weight decay was set to $5.10\text{-}4$. Training was conducted for 70 epochs with a mini-batch size of 16 on an NVIDIA A40 GPU using PyTorch [42].
A.5 Baseline CAM-based methods
Gradient-based methods primarily differ in how gradients are aggregated to compute importance weights, while gradient-free methods mainly vary in how the weights are computed.
ScoreCAM
[56]. A gradient-free method that eliminates the need for gradient information by assessing the importance of each activation map based on its forward-pass contribution to the target class score, and produces the final output via a weighted sum of these maps.
LayerCAM
[30]. A gradient-based method that generates class activation maps by leveraging the element-wise product of ReLU-activated gradients and feature maps at any convolutional layer, enabling fine-grained, spatially precise visual explanations without requiring global average pooling.
GradCAM
[47]. A gradient-based approach that uses the gradients of the target class flowing into the final convolutional layer to produce a coarse localization map, highlighting important regions in the image by upsampling the resulting map.
Guided backpropagation (Guided BP)
[51]. A gradient-based approach that modifies the standard backpropagation process to propagate only positive gradients through positive activations, producing fine-grained visualizations that highlight features strongly activating specific neurons in relation to the target output.
Integrated Gradient (Itgt Grad.)
[53]. A gradient-based method that attributes model predictions to input features by computing the path integral of gradients along a straight-line path from a baseline to the actual input, yielding fine-grained explanations.
Appendix B Explainability metrics
B.1 Activation consistency
The activation consistency [18] quantifies how well local explanations (e.g., individual activations within explanation maps) globally reflect the disease and healthy samples across a dataset. Specifically, it measures whether the activation patterns in the explanation maps consistently reflect the underlying disease or healthy class labels.
Following [18], we evaluated activation consistency by computing the proportion of positive activations (indicative of disease evidence) in saliency maps of disease samples, and negative activations (indicative of the absence of disease) in those of healthy samples. These proportions were calculated over the test set to assess whether the heatmaps consistently highlight pathological features in diseased cases and suppress activations in healthy ones. This metric thus captures the alignment between the semantic meaning of activations and the ground truth labels, offering a dataset-level evaluation of the coherence of local explanations with the global classification objective.
B.2 Activation precision and activation sensitivity
Let $\mathcal{X}=\{\bf{X}\}_{i=1}^{n}$ denote a set of input images, =$\mathcal{M}=\{\bf{M}\}_{i=1}^{n}$ the corresponding binary segmentation masks, and $\mathcal{S}=\{\bf{S}\}_{i=1}^{n}$ the associated explanation or saliency maps generated by any method. Activation precision measures the proportion of the saliency map’s positive mass that lies within the annotated region (the segmentation mask) [9, 43]. To compute it, saliency maps are first preprocessed by setting negative values to zero while retaining all positive values. This highlights how much of the explanation signal aligns with human-annotated ground truth, effectively quantifying the precision of an explainability method. The activation precision is defined as:
$AP(\mathcal{M},\mathcal{S})=\frac{\sum_{M,S}\sum_{i,j}M_{i,j}.S_{i,j}}{\sum_{i,j}S_{i,j}}.$ (6)
However, activation precision does not penalize false-negative (i.e. missed relevant regions). To address this, we introduce activation sensitivity, which captures the completeness of the explanation by evaluating how much of the annotated region is covered by the saliency map. The activation precision is defined as:
$AS(\mathcal{M},\mathcal{S})=\frac{\sum_{M,S}\sum_{i,j}M_{i,j}.S_{i,j}}{\sum_{i,j}M_{i,j}}.$ (7)
Unlike activation precision, activation sensitivity penalizes low saliency values within the mask. For example, if $M_{i,j}=1$ but $10<S_{i,j}<$ 1, the low activation will contribute little to the numerator, reflecting reduced confidence in that region. This makes activation sensitivity especially relevant in clinical tasks where completeness is critical, such as identifying multi-focal infectious diseases like pneumonia [37].
B.3 Top-k localization precison
Top-k localization precision [18] measures the ability of an explanation map to correctly highlight salient regions that overlap with ground-truth annotations. Specifically, it quantifies the proportion of the top-k positively activated regions within an explanation that match with annotated areas. In our implementation, each explanation map is first upsampled to the input resolution and then split into non-overlapping patches of size $33\times 3$ 33. These patches are ranked based on their average activation, and the top-k (k=30) most salient patches are selected. The precision is then computed as the fraction of these patches that overlap with the annotated ground-truth regions.
This metric can be viewed as a generalization of the pointing game metric [60], where only the single most activated region (top-1) is considered, to multiple regions, making it more suitable for medical imaging tasks. In such contexts, disease-relevant features (e.g., retinal lesions or pathological markers) are often spatially distributed across the image, rather than confined to a single localized area.
B.4 Failthfulness
Faithfulness, also referred to as sensitivity or fidelity [39], is a widely used metric to evaluate how accurately an explanation reflects the model’s true decision-making process. It assesses whether the importance scores (attributions) assigned to input features correspond to the actual impact of those features on the model’s prediction.
In our implementation, we focus on correctly classified samples from the test set. For each, the corresponding explanation map is upsampled to the input resolution and split into non-overlapping patches of size $33\times 3$ 33. These patches are ranked based on their mean activation values, and the top-k (k=30) most salient patches are iteratively occluded. After each occlusion step, we recorded the relative drop in the model’s confidence score for the predicted class. This process yields a deletion curve, from which we compute the Area Under the Deletion Curve (AUDC). A lower AUDC indicates a more faithful explanation, as it reflects a greater decline in model confidence when the most important regions (as indicated by the explanation map) are removed, suggesting that those regions were indeed critical to the model’s prediction.
Appendix C Effect of Lasso regularization on model performance for the binary tasks
The Lasso regularization coefficient $\lambda_{1}$ in Eq. 5 controls the sparsity of the class evidence map, encouraging the model to localize disease regions with high precision. For each task, $\lambda_{1}$ was selected based on a trade-off between accuracy and AUC on the corresponding validation set, choosing the highest values for which classification performance did not degrade significantly.
Figure 6: Model selection on validation sets under varying Lasso regularization strengths. The regularization coefficient $\lambda$ influences model performance, with notable effects on some datasets but minimal impact on the OCT dataset. The red markers indicate the selected $\lambda$ values, chosen to balance sparsity and classification performance.
Appendix D Additionnal Results
D.1 SoftCAM provides inherently interpretable visual explanations
Figure 7: Example explanations generated by different methods from VGG-16. The first column shows disease images with reference annotations, indicated by green markers or bounding boxes. Each row, from top to bottom, corresponds to fundus, OCT, and Chest X-ray images, respectively. The next five columns present saliency maps generated by post-hoc explanation methods, gradient-free (ScoreCAM, LayerCAM) and gradient-based (GradCAM, Guided BP, Itgd Grad). The final two columns showcase our proposed inherently interpretable dense and sparse SoftCAM explanations.
Figure 8: Additional example explanations of disease images from the ResNet model.
Figure 9: Additional example explanations of disease images from the VGG model.
D.2 Activation consistency
We quantify activation consistency only for the SoftCAM variants, as post-hoc methods are not inherently explainable, meaning their explanations do not directly influence the model’s decision-making process.
The results align well with qualitative visualizations. On the Fundus dataset, the sparse SoftCAM model exhibits a higher proportion of positive activations with the ResNet backbone, attributed to reduced false positives from the dense model, and fewer negative activations, reflecting the suppression of low-importance activations to zero. On the VGG backbone, regularization primarily reduces false-positive activations from the dense model but leads to a slight increase in activations on healthy samples. Similar result can be observed on the RSNA dataset.
On the OCT dataset, the dense SoftCAM with the ResNet backbone generally produces coarse-grained evidence around lesion areas. In contrast, the sparse variant refines these explanations, resulting in lower positive and negative activations across both disease and healthy samples, suggesting more selective and focused localization. However, with the VGG backbone, a higher proportion of negative activations is observed, reflecting the impact of the regularization strength, highlighting the importance of appropriately tuning this parameter for different architectures.
Table 2: Activation consistency on the ResNet model. $\mathbf{r_{LG}^{+}}$ denotes the proportion of positive or disease activations from disease images, while $\mathbf{r_{LG}^{-}}$ refers to the proportion of negative or healthy activations from healthy images.
Fundus OCT RSNA
$\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$ $\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$ $\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$
dense SoftCAM $0.28\pm 0.1$ $0.86\pm 0.1$ $0.30\pm 0.1$ $0.85\pm 0.1$ $0.75\pm 0.1$ $0.47\pm 0.1$
sparse SoftCAM $0.55\pm 0.2$ $0.76\pm 0.2$ $0.23\pm 0.1$ $0.83\pm 0.1$ $0.79\pm 0.1$ $0.45\pm 0.1$
Table 3: Activation consistency on the VGG model. $\mathbf{r_{LG}^{+}}$ denotes the proportion of positive or disease activations from disease images, while $\mathbf{r_{LG}^{-}}$ refers to the proportion of negative or healthy activations from healthy images.
Fundus OCT RSNA
$\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$ $\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$ $\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$
dense SoftCAM $0.32\pm 0.2$ $0.93\pm 0.1$ $0.75\pm 0.11$ $0.51\pm 0.1$ $0.75\pm 0.1$ $0.51\pm 0.1$
sparse SoftCAM $0.28\pm 0.2$ $0.94\pm 0.1$ $0.35\pm 0.14$ $0.95\pm 0.1$ $0.35\pm 0.1$ $0.95\pm 0.1$
Overall, the effect of regularization on the explanations varies depending on the backbone architecture. Nevertheless, the activation consistency metric aligns well with the qualitative explanations, generally capturing the impact of regularization across the dataset for a given architecture.
D.3 Precision and sensitivity analysis
We quantitatively evaluate the explanations generated by various methods using the ResNet and VGG backbones on the RSNA dataset. With the ResNet model, the dense SoftCAM achieves the highest localization precision, whereas the sparse SoftCAM yields the best results in terms of sensitivity. This discrepancy underscores the importance of developing evaluation metrics that balance human-aligned localization quality with model fidelity, capturing both interpretability and decision relevance.
Figure 10: Precision vs. sensitivity analysis on the RSNA dataset. Quantitative evaluation of explanations generated by different methods from the ResNet and VGG models on the RSNA dataset.
D.4 SoftCAM provides localized and faithful explanations
Table 4: Top-k localization precision and sensitivity. Sensitivity is quantified as the Area Under the Deleted Curve (AUDC), where lower values indicate greater faithfulness—that is, a larger drop in the model’s confidence when the most relevant patches are removed. For precision, higher values indicate better alignment between saliency maps and ground truth annotations. We refer to AUDC as “Del” and Top-K as “Top” with $K=30$.
ResNet (Topk $\uparrow$, AUDC $\downarrow$) VGG (Topk $\uparrow$, AUDC $\downarrow$)
Fundus OCT RSNA Fundus OCT RSNA
Top Del Top Del Top Del Top Del Top Del Top Del
ScoreCAM $0.16$ $0.67$ $0.07$ $0.73$ $0.66$ $0.97$ $0.20$ $0.67$ $0.08$ $0.55$ $0.62$ $0.88$
LayerCAM $0.22$ $0.65$ $0.08$ $0.74$ $0.65$ $0.97$ $0.23$ $0.64$ $0.08$ $0.56$ $\bf{0.65}$ $0.84$
GradCAM $0.37$ $0.64$ $0.14$ $0.73$ $0.67$ $0.95$ $0.65$ $0.68$ $0.14$ $0.58$ $0.61$ $0.86$
Guided BP $0.30$ $\bf{0.57}$ $0.23$ $0.68$ $0.55$ $0.97$ $0.43$ $0.57$ $0.23$ $\bf{0.40}$ $0.58$ $0.85$
Itgd Grad. $0.28$ $0.63$ $0.20$ $0.70$ $0.52$ $0.98$ $0.33$ $\bf{0.62}$ $0.2$ $0.51$ $0.55$ $0.88$
dense SoftCAM $0.39$ $0.69$ $0.46$ $0.61$ $0.65$ $\bf{0.92}$ $0.54$ $0.63$ $0.72$ $0.45$ $0.64$ $0.84$
sparse SoftCAM $\bf{0.52}$ $0.68$ $\bf{0.86}$ $\bf{0.31}$ $\bf{0.73}$ $0.93$ $\bf{0.65}$ $0.674$ $\bf{0.82}$ $0.43$ $0.63$ $\bf{0.82}$
Appendix E Activation precision and sensitivity on the RSNA dataset
E.1 Lasso vs Ridge penalty
Figure 11: Model selection on validation sets under varying Lasso and Ridge regularization strengths. The regularization coefficients $\lambda_{1}$ and $\lambda_{2}$ influence model performance. The red markers indicate the selected regularization values, chosen to balance classification performance.
E.2 Activation precision vs. activation sensitivity
Table 5: Activation Precision (AP) vs. Activation Sensitivity (AS) for different SoftCAM variants and baseline post-hoc methods. The dense SoftCAM consistently lies between the lasso and ridge variants, highlighting the importance of balancing the two regularization terms to achieve an optimal trade-off between precision and completeness in the explanations.
ResNet VGG
AP $\uparrow$ AS $\uparrow$ AP $\uparrow$ AS $\uparrow$
ScoreCAM $0.470$ $0.318$ $0.403$ $0.303$
LayerCAM $0.456$ $0.300$ $0.401$ $0.120$
GradCAM $0.525$ $0.252$ $0.373$ $0.260$
Guided BP $0.381$ $0.033$ $0.364$ $0.044$
Itgd Grad. $0.286$ $0.040$ $0.322$ $0.039$
dense SoftCAM $0.526$ $0.251$ $0.461$ $0.355$
sparse SoftCAM $\bf{0.654}$ $0.182$ $\bf{0.519}$ $0.320$
lasso SoftCAM $0.440$ $\bf{0.316}$ $0.412$ $\bf{0.396}$
E.3 More examples: activation precision vs activation sensitivity
Figure 12: Additional examples of localization evaluation on the RSNA dataset for pneumonia detection. Each column shows explanation maps generated by different methods. Ground-truth bounding boxes are overlaid on each map, with the top-right value indicating the activation precision, while the top-left value indicates the activation sensitivity. The high precision from the lasso model and the more complete explanations from the ridge model emphasize the importance of balancing the two regularization terms to achieve an optimal trade-off.
Appendix F Multi-class analysis
F.1 Regularization
Given the small size of retinal lesions, we used Lasso regularization, selecting $\lambda_{1}$ values that balanced performance (Fig. 13)
Figure 13: Model selection on validation sets under varying Lasso regularization strengths. The regularization coefficients $\lambda_{1}$ influence model performance. The red markers indicate the selected regularization values to balance classification performance.
F.2 Faithfulness
As no ground-truth lesion annotations were available for the multi-class tasks, we evaluated the faithfulness of the explanations by measuring their contribution to model predictions. For correctly classified test samples, we progressively removed top-k ($k=30$) ranked patches (based on the explanation maps) and tracked the average drop in class confidence (Fig. 14).
Figure 14: Sensitivity analysis.
F.3 Area Under the Deleted Curve
The area under the deletion curve (AUDC) was computed from the sensitivity analysis (Fig. 14). In both tasks, the dense and sparse SoftCAM achieved superior performance, with sparse SoftCAM yielding the lowest AUDC, indicating the highest faithfulness (Tab. 6).
Table 6: Area Under the Deleted Curve (AUDC $\downarrow$).
ResNet VGG
Fundus OCT Fundus OCT
ScoreCAM $0.894$ $0.819$ $0.880$ $0.852$
LayerCAM $0.889$ $0.817$ $\bf{0.869}$ $0.850$
GradCAM $0.887$ $0.815$ $0.872$ $0.847$
Guided BP $0.899$ $0.793$ $0.905$ $0.823$
Itgd Grad. $0.907$ $0.821$ $0.901$ $0.833$
dense SoftCAM $0.870$ $0.825$ $0.905$ $0.826$
sparse SoftCAM $\bf{0.856}$ $\bf{0.609}$ $0.882$ $\bf{0.806}$
F.4 Qualitative explanation on retinal fundus images
For the multi-class tasks on DR detection from fundus images, SoftCAM variants produced more focused and class-consistent explanations. In addition to the sparse and dense evidence maps, we also provide visualizations for post-hoc methods.
Figure 15: Class-specific explanation with the ResNet backbone. The application of our method to multi-class DR detection demonstrates the utility of class-specific explanations produced by the sparse SoftCAM, which more precisely highlight disease-relevant regions compared to the dense SoftCAM and the best-performing post-hoc method, GradCAM. In the example shown, the image is labeled as severe DR, and the highlighted regions correspond to suspicious areas, reflecting relevant DR lesions.
Figure 16: Class-specific explanation with the VGG backbone. The application of our method to multi-class DR detection demonstrates the utility of class-specific explanations produced by the sparse SoftCAM, which more precisely highlight disease-relevant regions compared to the dense SoftCAM and the best-performing post-hoc method, ScoreCAM. In the example shown, the image is labeled as severe DR, and the highlighted regions correspond to suspicious areas, reflecting relevant DR lesions.
F.5 Qualitative explanation on retinal OCT images
For the multi-class tasks on retinal disease classification from OCT images, SoftCAM variants produced more focused and class-consistent explanations. In addition to the sparse and dense evidence maps, we also provide visualizations for GradCAM and Guided BP, as these were the best-performing post-hoc methods for the ResNet and VGG backbones, in terms of the Area Under the Deletion Curve (Tab. 6).
Figure 17: Class-specific explanation with the ResNet backbone. SoftCAM applied to multi-class retinal disease classification demonstrates the utility of class-specific explanations, with the sparse SoftCAM variant more precisely highlighting disease-relevant regions compared to both the dense SoftCAM and the best-performing post-hoc methods, GradCAM and Guided Backpropagation. In the example shown, the image is labeled as Diabetic Macular Edema (DME), and the highlighted regions produced by sparse SoftCAM highlight suspicious areas, reflecting relevant retinal lesions.
Figure 18: Class-specific explanation with the VGG backbone. SoftCAM applied to multi-class retinal disease classification demonstrates the utility of class-specific explanations, with the sparse SoftCAM variant more precisely highlighting disease-relevant regions compared to both the dense SoftCAM and the best-performing post-hoc methods, GradCAM and Guided Backpropagation. In the example shown, the image is labeled as Diabetic Macular Edema (DME), and the highlighted regions produced by sparse SoftCAM highlight suspicious areas, reflecting relevant retinal lesions.

### v2

@@
-Soft-CAM: Making black box models self-explainable for high-stakes decisions
+
 Abstract
@@
 By removing the global average pooling layer and replacing the fully connected classification layer with a convolution-based class evidence layer, SoftCAM preserves spatial information and produces explicit class activation maps that form the basis of the model’s predictions.
-Evaluated on three medical datasets, SoftCAM maintains classification performance while significantly improving both the qualitative and quantitative explanation compared to existing post-hoc methods.
-Our results demonstrate that CNNs can be inherently interpretable without compromising performance, advancing the development of self-explainable deep learning for high-stakes decision-making.
+Evaluated on three medical datasets spanning three imaging modalities, SoftCAM maintains classification performance while significantly improving both the qualitative and quantitative explanation compared to existing post-hoc methods.
+The code is available at https://github.com/kdjoumessi/SoftCAM.
+keywords:
+ElasticNet, Self-explainable models, Convolutional Neural Networks, Class Activation Maps, Attribution maps.
+editors: Under Review for MIDL 2026
 1 Introduction
-Convolutional Neural Networks (CNNs) have revolutionized computer vision by efficiently capturing local patterns, reducing parameters, and accelerating convergence, enabling superior performance in tasks like image recognition and object detection [35, 40].
-However, their lack of interpretability limits adoption in high-stakes fields like medicine, where transparency and trust are crucial.
-To explain CNNs, numerous saliency-based methods have been proposed, including class activation maps (CAM) [61] and their variants [14, 47, 56, 59], gradient-based techniques [48, 50, 51, 53], and even perturbation- or occlusion-based methods [22, 56, 59].
-These methods have been widely adopted to explain the decisions of trained black-box models.
-Such saliency map-based techniques offer explanations for CNN classifiers that claim to highlight regions in the input image most relevant to the model’s prediction.
-These explanations are generated post-hoc, typically after a model is trained [10, 27].
-Studies have shown significant limitations in their effectiveness, especially in clinical settings [5].
-Post-hoc saliency methods often lack faithfulness, reliability, and consistency, resulting in explanations that may not accurately reflect the model’s decision-making process [3, 58].
-Moreover, they struggle to localize relevant regions in medical imaging [4], where the limited availability of ground-truth annotations makes it difficult to assess their trustworthiness.
-To overcome these challenges, inherently and/or self-explainable models have been introduced [45], designed explicitly to provide interpretable insights by incorporating explanations within their architecture [11, 13, 16, 18, 52].
-These models generate more trustworthy and faithful explanations that align closely with the model’s actual reasoning [28].
-However, self-explainable models generally use specific architectures [13, 16, 52], which limits their applicability and generalization to widely used CNN architectures.
-Motivated by these challenges, we propose SoftCAM, a straightforward generalization of Class Activation Maps (CAM) that uses a convolution-based classifier to transform any black-box CNN into a self-explainable model.
-By removing the final pooling layer and replacing the fully connected classification layer with 1x1 convolutions, SoftCAM turns classical CNNs into fully convolutional networks, generating class-specific evidence maps that are directly used for predictions.
-Our contributions are:
-  • We introduced SoftCAM, a simple modification to CNNs that enables self-explainability, and experimentally demonstrate that the resulting model preserves classification performance across three clinically relevant medical datasets spanning different imaging modalities.
-  • We showed that regularizing evidence maps using ElasticNet, a regularizer combining both ridge and lasso penalties, enhances the model’s explanations.
-  • We evaluated five widely used traditional CAM-based post-hoc explanation methods, showing that SoftCAM most often outperforms them across a broad range of explainability metrics and across all three considered medical imaging datasets and modalities.
-2 Related work
-Deep neural networks (DNNs) are widely used in a variety of fields.
-However, regulatory frameworks such as the European AI Act require AI-based decisions to be explainable to ensure fairness, transparency, and accountability, allowing users to understand their decision-making process [2, 41].
-Explainability can help verify and even improve performance by detecting shortcuts and identifying clinically relevant features, ultimately fostering greater trust in decision-making, especially in fields like healthcare [17].
-Existing explainable AI (XAI) methods for image analysis can be broadly categorized into attribution-based and non-attribution-based approaches [10, 27, 57].
-Attribution-based methods explain “where” important features exist in the input by generating saliency or heatmaps that assign importance scores to individual pixels or regions, helping visualize their contribution to the model’s decision.
-These include perturbation-based methods [29] and class activation maps [26], which consist of both gradient-based [47, 51, 53] and gradient-free approaches [44, 56, 61].
-In contrast, non-attribution-based approaches explain “why” a decision was made without relying on importance scores, instead using techniques such as concept-based (ACE [23], TCAV [32], CBM [33]), prototype-based [15, 16], or counterfactual-based [7, 12, 24, 52] methods to analyze model behavior from different perspectives.
-These approaches also differ in how explanations are obtained and which architectures they can be applied to.
-Attribution-based methods typically provide post-hoc, local explanations by offering input-specific insights into black-box CNN models after training.
-In contrast, non-attribution-based methods are generally inherently interpretable by design, promoting transparency and enabling global understanding of the model’s decision-making process across the entire dataset [45].
-However, in some cases, self-explainable models may be less effective for complex tasks, highlighting a tradeoff between interpretability and performance, where increasing transparency may sometimes come at the cost of classification performance [57].
-Our method, SoftCAM, relies on an explicit class-evidence layer based on convolutional operations for classification, offering class-specific, attribution-based explanations while remaining inherently interpretable, unlike post-hoc approaches.
-Moreover, it maintains predictive performance comparable to its corresponding non-interpretable black-box models.
-Closely related work includes [6] and [18].
-In [6], a dual-branch approach is used, where one branch leverages a traditional black-box CNN model (ResNet-18) for classification, and the second branch uses weight-sharing for post-hoc explanations, requiring two forward passes for inference.
-In the second branch for post-hoc explanation, the global average pooling layer (GAP) is removed, and the linear classifier is replaced by convolutional layers that share weights during inference to generate class-specific activation maps.
-In contrast, SoftCAM is trained end-to-end, providing both predictions and explanations in a single forward pass, eliminating the need for additional computational overhead or weight sharing.
-Furthermore, while [18] uses explicit class-evidence maps to enhance the explanation of a self-explainable bag-of-local-feature model (BagNet [13]), our method transforms black-box models into self-explainable models.
-We evaluated our approach on a range of medical datasets, comparing the resulting explanations to various post-hoc attribution-based methods, including both gradient-free and gradient-based techniques.
+Convolutional Neural Networks (CNNs) have revolutionized computer vision by effectively capturing local spatial patterns, reducing the number of parameters, and enabling faster convergence, leading to state-of-the-art performance in tasks such as image recognition and object detection [oquab2015object, li2021survey].
+However, their lack of interpretability limits adoption in high-stakes fields such as medical image analysis, where transparency and trust are essential.
+To mitigate this issue, numerous post-hoc saliency- or attribution-based methods have been proposed to explain CNN predictions by highlighting “where” important features appear in the input through pixel-wise importance maps computed after training.
+Prominent approaches include class activation maps (CAM) techniques [zhou2016learning, selvaraju2017grad, jiang2021layercam], which combine convolutional features with class-specific weights; backpropagation-based methods [springenberg2014striving, sundararajan2017axiomatic], which propagate output gradients to the input to estimate pixel relevance; and perturbation-based methods [wang2020score, ivanovs2021perturbation], which systematically alter input features and measure the corresponding change in the model’s output.
+While post-hoc attribution-based methods provide intuitive visualizations of discriminative regions, they frequently approximate rather than accurately reflect the model’s true reasoning [adebayo2018sanity, arun2021assessing, saporta2022benchmarking] and are often valid only under strong regularity assumptions [gunther2025informative].
+Their post-hoc nature further limits their effectiveness—particularly in clinical applications [arun2021assessing]—due to low faithfulness, reliability, and consistency.
+Consequently, their visual outputs may not accurately reflect the model’s internal decision-making [adebayo2018sanity, saporta2022benchmarking].
+Moreover, post-hoc techniques often struggle to precisely localize disease-relevant regions in medical images [arun2021assessing], where the scarcity of ground-truth annotations further complicates validation.
+To address these challenges, inherently or self-explainable models have been proposed [rudin2019stop], embedding interpretability directly into their architectures [brendel2018bagnets, chen2019looks, koh2020concept].
+By coupling prediction with explanation, self-explainable models provide interpretable and transparent insights, in line with human reasoning.
+However, their reliance on specialized architectures limits generalization to widely used CNN models.
+Motivated by these limitations, we introduce SoftCAM, a simple yet effective framework that makes CNNs inherently interpretable without relying on post-hoc explanation methods.
+SoftCAM generalizes the concept of class activation maps to transform conventional CNNs into self-explainable models.
+By removing the final global average pooling layer and replacing the fully connected classifier with a convolution-based class-evidence layer, SoftCAM converts standard CNNs into fully convolutional architectures that generate explicit, class-specific evidence maps used both for prediction and visual explanation.
+Evaluated on two widely used CNN architectures, we showed that the resulting SoftCAM-based models maintain competitive accuracy with their black-box baselines while achieving superior interpretability across three clinically relevant medical imaging datasets spanning three modalities.
+Furthermore, applying ElasticNet regularization, which combines ridge and lasso penalties, on evidence maps improves explanations both qualitatively and quantitatively, while revealing a task-dependent trade-off between sparsity and density.
+Finally, a comprehensive evaluation against five widely used post-hoc attribution methods showed that SoftCAM consistently outperforms these approaches on various explainability metrics.
 2 Method
 Preliminaries
-Given an input image $\mathbf{X}\in\mathbb{R}^{H_{X}\times W_{X}\times C_{X}}$ with height $H_{X}$, width $W_{X}$, and the number of channels $C_{X}$, consider a CNN network $f_{\theta}$ that maps $\mathbf{X}$ to a probability distribution $\mathbf{\hat{y}}=f_{\theta}(\mathbf{X})\in\mathbb{R}^{C}$, where $C$ is the number of classes, and $y^{c}\in\mathbf{y}$ represents the predicted probability for class $c$.
-The network consists of a feature extractor $g_{\phi}$, and a classifier layer $h_{\psi}$, with learnable parameters $\phi$ and $\psi$.
-The feature extractor generates a feature map $\mathbf{Z}=g_{\phi}(\mathbf{X})\in\mathbb{R}^{N\times M\times D}$, where $N\times M$ denotes the spatial size and $D$ is the feature dimension (e.g., $D=2048$ for most ResNet variants).
-The classifier then predicts the final output based on the extracted features.
-Let $\mathcal{A}=\{\mathbf{A}_{k}\}_{k=1}^{D}$, the set of activation maps obtained from the feature extractor, where $A_{k}$ is the activation of the $k$-th neuron.
-Let, $S^{c}_{\text{Map}}\in\mathbb{R}^{N\times M}$ be the 2D saliency map, providing a visual explanation of the model’s prediction for class $c$.
-This paper explores how to train self-explainable CNNs to simultaneously generate both the prediction $y^{c}$ and its corresponding explanation $S^{c}_{\text{Map}}$.
-Traditional CNN architectures employ a GAP layer to reduce the feature map to $1\times D$, followed by a classification module consisting of one or more linear fully connected layers (FCLs) to generate the final prediction.
-Post-hoc methods are then typically used to explain the model’s decision.
+[…]
+In contrast, traditional CNN are black-boxes: a global average pooling layer reduces $\mathbf{Z}$ to a vector of size $1\times D$, which is then fed into one or more linear fully connected layers (FCL) to generate the final prediction.
+Post-hoc methods are required to explain decisions.
 2.1 CAM-based methods
-Class Activation Maps (CAMs) [61] are closely related to our approach, offering local visual explanations of CNN predictions by generating saliency maps for individual inputs.
-CAM achieves this by linearly combining the feature maps from the final convolutional layers with importance coefficients from the FCL classifier, thereby producing class-wise attribution maps as follows:
-$S_{\text{CAM}}^{c}(x_{1},x_{2})=\sum_{k=1}^{D}w_{k}^{c}A_{k}(x_{1},x_{2}),$ (1)
-where $A_{k}(x_{1},x_{2})$ is the activation of neuron $k$ in the feature map at spatial location $(x_{1},x_{2})$, and $w_{k}^{c}$ denotes the importance weight associated with class $c$ for unit $k$ in the fully connected layer.
-Originally, CAM was designed for CNNs with GAP and FCL, but has been extended to gradient-based methods using class score gradients to compute importance weights [14, 47, 48].
-This extension enabled CAM-based techniques to be applied to a broader range of CNN architectures, particularly those where the GAP layer is followed by multiple FCLs, as seen in models like VGG [49] and InceptionV3 [54].
-For example, GradCAM [47] extends the original CAM approach by backpropagating the gradient from a target class to the input layer to highlight the image regions that strongly influence the model’s prediction.
-GradCAM is formulated as
-$S^{c}_{\text{Grad-CAM}}(x_{1},x_{2})=\text{ReLU}\Bigg{(}\sum_{k=1}^{D}w_{k}^{c}A_{k}(x_{1},x_{2})\Bigg{)},$ (2)
-where the weight coefficients are computed as $w_{k}^{c}=\frac{1}{N\times M}\sum_{i}^{N}\sum_{j}^{N}\frac{\partial y^{c}}{\partial A_{k}(i,j)}$.
-Here, $A_{k}(i,j)$ is the activation value at location $(i,j)$ on $A_{k}$, and the rectified linear unit (ReLU) is applied to ensure that the final activation map considers only the features that positively influence class $c$.
-Following GradCAM, several variations have been proposed, including gradient-based approaches such as SmoothGrad [50], GradCAM++ [14], guided-backpropagation [51], and integrated gradients [53], as well as gradient-free methods like ScoreCAM [56], LayerCAM [30], and OptiCAM [59].
-Gradient-based methods primarily differ in how gradients are aggregated to compute importance weights, while gradient-free methods mainly vary in how the weights are computed.
-Despite the success of class activation map-based methods in explaining CNN classifiers, including medical applications [8], they have a key limitation: they rely on already trained models and provide post-hoc explanations, which may not accurately reflect the model’s true decision-making process.
-Additionally, gradient-based methods face inherent challenges such as gradient saturation, where DNN gradients tend to diminish, and false confidence, where the highest activation map weight does not necessarily correspond to the greatest increase in confidence [56].
-On the other hand, gradient-free methods are computationally and memory-intensive, as they often require multiple forward passes on perturbed inputs.
-Finally, CAM-based methods are easy to implement in CNNs with clearly defined spatial feature maps, but face challenges in multi-branch architectures like InceptionV3 due to the complexity of integrating diverse feature maps from parallel convolutional paths.
+Class Activation Maps (CAM) [zhou2016learning] are closely related to our approach, providing visual explanations of CNN predictions through class-specific saliency maps.
+CAM operates by linearly combining the final convolutional feature maps with their corresponding importance weights from the FCL classifier to produce class-wise attribution maps:
+[…]
+where $A_{k}(x_{1},x_{2})$ denotes the activation of the $k$ feature map at location $(x_{1},x_{2})$, and $w_{k}^{c}$ is the class-specific importance weight for the $k$ feature map from the fully connected layer.
+Originally, CAM was designed for CNNs that use a GAP followed by a single FCL.
+Subsequent extensions, such as GradCAM [selvaraju2017grad] and LayerCAM [jiang2021layercam], introduced gradient-based approaches that use gradient to compute importance weights, enabling class-specific explanations for various architectures—particularly those with multiple FCLs after the GAP layer, such as VGG [simonyan2014very].
+In GradCAM, the importance weights are computed by globally averaging the gradients of the target class score as $w_{k}^{c}=\frac{1}{N\times M}\sum_{i}^{N}\sum_{j}^{N}\frac{\partial y^{c}}{\partial A_{k}(i,j)}$, where $A_{k}(i,j)$ represents the activation at spatial location $(i,j)$ in the $k$-th feature map.
+Following GradCAM, several gradient-based and gradient-free extensions have emerged [he2022survey].
+While gradient-based methods differ primarily in how they aggregate gradients to compute importance weights, gradient-free approaches estimate these weights without backpropagation, often relying on perturbation-based methods like ScoreCAM [wang2020score].
+Despite their widespread clinical use [ayhan2022clinical], attribution-based methods have notable limitations [gunther2025informative]: they provide post-hoc explanations that may not reflect the model’s true reasoning (Appendix A).
+Gradient-based variants can suffer from gradient saturation and false confidence [wang2020score], while gradient-free approaches are computationally costly, requiring many forward passes on perturbed inputs.
+2.2 Generalizing CAMs for self-explanability
+Motivated by the limitations of post-hoc class activation map-based methods in interpreting CNN, we introduce SoftCAM (Fig.
+1), a straightforward modification of black-box CNN classifiers that makes them inherently interpretable.
+SoftCAM achieves this by replacing the fully connected classification layer in classical CNNs with an explicit class-evidence convolutional layer, preserving spatial information and providing explanations in a single forward pass, eliminating the need and computational overhead for post-hoc techniques.
 Figure 1: Overview of softCAM for making black-box CNNs inherently interpretable.
@@
 Class-specific evidence maps (g) are upsampled and overlaid on the input to visualize the model’s decision-making process.
-3.2 Improving CAMs for self-explanability
-Motivated by the limitations of post-hoc class activation map-based methods in interpreting CNN classifiers, we introduce SoftCAM (Fig.
-1), a straightforward modification of black-box CNN classifiers that makes them self-explainable and inherently interpretable.
-SoftCAM achieves this by replacing the fully connected classification layer in classical CNNs with an explicit class-evidence convolutional layer, preserving spatial information and providing explanations in a single forward pass, eliminating the need and computational overhead for post-hoc techniques.
-We make black-box CNN architectures self-explainable by modifying how predictions are obtained.
-Any FCL of size $b_{1}\times b_{2}$, where $b_{1}$ and $b_{2}$ denote the number of input and output features, respectively, can be equivalently expressed as a $1\times 1$ convolutional layer with $b_{1}$ input channels and $b_{2}$ output channels [18].
-This allows us to replace FCL classifiers in standard CNN architectures with convolutions, removing the GAP layer before classification, while preserving model complexity and spatial localization.
+We make black-box CNN self-explainable by modifying how predictions are obtained.
+Any FCL of size $b_{1}\times b_{2}$ can be equivalently expressed as a $1\times 1$ convolutional layer with $b_{1}$ input channels and $b_{2}$ output channels.
+This allows systematic replacement of FCL classifier heads in CNN classifiers with convolutional layers, eliminating the GAP layer before classification while preserving model complexity and spatial localization.
 The new classifier module $h$ consists of convolutional layers (Fig.
-1d) with $C$ convolution kernels of size $1\times 1$ and unit stride, producing class evidence maps (Fig.
-1e)
+1d) with $C$ convolution kernels of size $1\times 1$ and unit stride, producing class-specific evidence maps (Fig.
+1e):
 $\mathbf{A}=h_{\psi}(\mathbf{Z})\in\mathbb{R}^{M\times N\times C},$ (2)
 where $\psi$ is a learnable parameter.
-Indeed, $h_{\psi}$ can be viewed as an explainable, soft generalization of classical post-hoc attribution methods (Eq.
-1, 2), mapping the low-dimensional feature volume $\mathbf{Z}$ into an interpretable, class-wise activation volume $\mathbf{A}$ whose reduced channel dimension corresponds to the number of target classes.
+These maps can be upsampled to the input resolution and overlaid on the image (Fig.
+1g) for visualization and clinical interpretation.
+The module $h_{\psi}$ can be viewed as a soft, explainable generalization of classical post-hoc attribution methods (Eq.
+1), mapping the high-dimensional feature map $\mathbf{Z}$ with $D$ channels into low-dimentional, class-wise activation maps $\mathbf{A}$ with $C$ channels corresponding to the target classes.
 Unlike CAM (Eq.
-1) and GradCAM (Eq.
-2), which generate post-hoc heuristic explanations, our approach leverages the feature map volume from the backbone and applies a parameterized function $h_{\psi}$ to directly produce class activation maps that are used for prediction.
+1) and related attribution methods which generate post-hoc heuristic explanations, our approach leverages the final feature map from the backbone and applies a parameterized function $h_{\psi}$ to directly produce class activation maps that are used for prediction.
 In contrast to classical CAM-based methods, the importance weights are not explicitly defined but are implicitly learned and encoded within the classifier’s parameters.
-The resulting architecture is a fully convolutional, self-explainable model, where the final predicted probabilities are computed from the evidence map (Fig.1e), without introducing additional parameters:
-$\mathbf{\hat{y}}=\text{Softmax}\bigg{(}\mbox{AvgPool}\Big{(}h_{\psi}\big{(}g_{\Phi}(\mathbf{X})\big{)}\Big{)}\bigg{)}\in\mathbb{R}^{1\times C}.$ (4)
-Additionally, the class evidence maps $\mathbf{A}$ serve as built-in explanations, directly representing the contribution of individual input regions to the final prediction (Fig.
+The resulting architecture is a fully convolutional, self-explainable model, where the predicted probabilities are derived directly from the evidence maps (Fig.1e), maintaining the model complexity without adding extra parameters:
+[…]
+Furthermore, the class evidence maps $\mathbf{A}$ serve as built-in explanations, directly representing the contribution of individual input regions to the final prediction (Fig.
 1g).
-Replacing linear FCL layers with convolutional operations offers several advantages.
-Due to the shift-invariance and position-agnostic properties of CNNs, all image regions are weighted equally when forming the final classification (Fig.
+Using class-evidence maps for classification offers several advantages: all importance scores are weighted equally when computing class probabilities (Fig.
 1f).
-As a result, input feature patches with high activations in the evidence maps contribute most significantly and linearly to the prediction.
-This behavior mirrors that of simple linear models, where each value in the activation map has a direct and interpretable impact on the output.
+Consequently, input feature patches with high activations in the evidence maps contribute more strongly to the prediction, analogous to linear models, where each input feature contributes linearly to the output.
 2.3 Regularizing SoftCAM for interpretability
-By using explicit class-evidence maps, the model can be trained directly with regularization applied to these maps to enhance interpretability.
-In practice, we apply an ElasticNet regularization [62], which linearly combines the $\ell_{1}$ (Lasso) and $\ell_{2}$ (Ridge) penalties, leading to the following loss function:
-$\mathcal{L}(\mathbf{y},\mathbf{\hat{y}})=\mbox{CE}(\mathbf{y},\mathbf{\hat{y}})+\lambda_{1}\sum_{i,j,c}|\mathbf{A}_{c}^{ij}|+\lambda_{2}\sum_{i,j,c}||\mathbf{A}_{c}^{ij}||_{2}.$ (5)
-Here, CE denotes the cross-entropy loss, and $\mathbf{y}$ represents the reference class labels.
-When $\lambda_{2}=0$, the ElasticNet penalty becomes the Lasso penalty, which promotes sparsity in the class evidence maps [18] by removing less informative activations, making it particularly useful for tasks where precision in explanations is crucial.
-In contrast, when $\lambda_{1}=0$, the ElasticNet reduces to the Ridge penalty, which reduces irrelevant activations without forcing them to zero, which is beneficial when minimizing false negatives is a priority.
-ElasticNet thus provides a balance between Lasso and Ridge penalties, balancing sparsity and smoothness in the resulting activation maps.
-Visualizing explanations.
-The evidence map generated by SoftCAM is upsampled to the input resolution for visualization (Fig.
-1g).
-Like most CAM-based methods, such as GradCAM [47], ScoreCAM [56], and LayerCAM [30] that operate on the final convolutional layer, SoftCAM’s explanations are limited by the resolution of the backbone (e.g., 16×16 for VGG-16/ResNet-50 with 512×512 input) due to pooling and striding, leading to lower-resolution saliency maps.
-However, by introducing the class evidence and classification layer directly atop features, SoftCAM regularizes the evidence map, making it less coarse and thereby enhancing localization.
-In contrast, gradient-based methods like Integrated Gradients [53] and Guided Backpropagation [51] produce high-resolution saliency maps by computing pixel-level gradients, which may lead to noisy maps, especially when the region of interest spans a broader area, as commonly observed in Chest X-ray images.
-Comparison with other approaches.
-Unlike post-hoc attribution-based approaches, our method is inherently interpretable from the classification layer and maintains performance comparable to its black-box counterpart, without a significant trade-off, even when regularization is applied to enhance explainability.
-Compared to [18], our method extends from the concept of interpretable bag-of-local models to general black-box CNN architectures and generalizes the regularization from Lasso to ElasticNet, with extensive evaluations across multiple datasets using a broad range of explanability metrics.
-Compared to [6], our method is trained end-to-end and does not require post-hoc processing, weight sharing between branches, or an additional forward pass to generate explanations.
+By using explicit class-evidence maps, SoftCAM-based models can be trained with regularization constraints directly applied to the explanation maps, enhancing interpretability.
+In practice, we apply an ElasticNet penalty [zou2005regularization], which linearly combines the $\ell_{1}$ (lasso) and $\ell_{2}$ (ridge) regularization, leading to the following loss function:
+[…]
+Here, CE denotes the cross-entropy loss, and $\mathbf{y}$ represents the reference labels.
+Setting $\lambda_{2}=0$, results in the lasso penalty promoting sparsity in evidence maps [donteu2023sparse] by suppressing less informative activations (mainly false positives), making it particularly useful for tasks where precision in explanations is crucial.
+In contrast, $\lambda_{1}=0$, gives the ridge penalty, which smooths activations without forcing them to zero, useful for assessing localization sensitivity over large regions, as it penalizes false negatives (Appendix E).
+ElasticNet thus provides a balance between lasso ($\lambda_{2}=0$) and ridge ($\lambda_{1}=0$) penalties.
 3 Experimental setup
 Datasets.
-We evaluated our approach on three publicly available medical datasets spanning three imaging modalities: the Kaggle Diabetic Retinopathy (DR) [20], Retinal OCT [31], and the RSNA Chest X-Ray (CXR) [1].
-The first dataset comprised high-resolution retinal color fundus images, each labeled with a DR severity score ranging from 00 (No DR) to $4$ (Proliferative DR).
-The second dataset included retinal OCT B-scans images categorized into Drusen, Diabetic macular edema, Choroidal neovascularization, and Normal cases.
-The final dataset consisted of high-resolution frontal-view chest radiographs labeled for pneumonia detection, with bounding box annotations for pneumonia cases.
-Additionally, lesion annotations were obtained for $65$ DR images from the Kaggle dataset [17] and $40$ Drusen images from the retinal OCT dataset [16].
-Each dataset was split into training, validation, and test sets using different dataset-specific train-validation-test proportions, ensuring that all samples from the same patient were assigned to the same split to prevent data leakage.
-For full details, see Appendix A.1.
+We evaluated our approach on three publicly available medical datasets spanning three imaging modalities: the Kaggle Diabetic Retinopathy (DR) [kaggle_dr_detection], Retinal OCT [kermany2018identifying], and the RSNA Chest X-Ray (CXR) [rsna_dataset].
+The first dataset comprised high-resolution retinal color fundus images labeled with DR severity score ranging from 0 (No DR) to $4$ (Proliferative DR).
+The second dataset included retinal OCT B-scans images labeled for three retinal disease conditions.
+The final dataset consisted of high-resolution frontal-view chest radiographs labeled for pneumonia detection, with bounding boxes for pneumonia cases.
+Additionally, clinicians annotated lesions in $65$ DR images from the Kaggle dataset [djoumessi2024inherently] and $40$ drusen images from the retinal OCT dataset [djoumessi2024actually] to allow directly clinical evaluation.
+Each dataset was split into training, validation, and test sets, ensuring that all samples from a given patient remained in the same split.
+For full details, see Appendix B.
 Baseline models.
-The effectiveness of our method was evaluated using two widely used black-box CNN architectures: ResNet-50 [25] and VGG-16 [49].
-They differ primarily in the design of their classification heads, where ResNet employs a single fully connected layer, while VGG uses multiple.
-In both models, we explicitly replaced the classification head with our convolutional evidence map layer, adapting the architecture to enable interpretability (see Appendix A.2).
-The models were sourced from Torchvision [36], initialized with pre-trained weights from ImageNet, and fine-tuned using a consistent setup [^1] Our code with datasets is available at https://anonymous.4open.science/r/SoftCAM-E1A3/.
-Training was performed over $70$ epochs with a mini-batch size of $16$ on an NVIDIA A40 GPU using PyTorch [42].
-A range of data augmentation and preprocessing techniques was applied (see Appendix A.3).
-For complete training details, see Appendix A.4.
-Baseline CAM-based methods.
-We qualitatively and quantitatively assessed the explanations generated by our method (SoftCAM) against post-hoc explanation techniques from several state-of-the-art class attribution map-based methods, applied to their respective black-box models.
-Specifically, we compared our approach with gradient-based methods, including GradCAM [47], Integrated Gradient (Itgd Grad.) [53], Guided Backpropagation (Guided BP) [51], as well as gradient-free methods such as ScoreCAM [56] and LayerCAM ([30]).
-Guided BP and Itgd Grad.
-have consistently performed well in producing saliency maps for explaining black-box CNN classifiers on retinal images [8, 17], while GradCAM has shown strong localization performance for chest X-ray interpretation [46].
-Gradient-based methods were implemented from Captum [34], whereas gradient-free methods were implemented via TorchCAM [21].
-For full descriptions of these methods, see Appendix A.5.
+Our method was evaluated using two widely used black-box CNN architectures: ResNet-50 [he2016deep] and VGG-16 [simonyan2014very].
+They differ primarily in the design of their classification heads, where ResNet employs a single fully connected layer, while VGG uses multiple layers.
+More details on the training setup2, including data preprocessing and augmentation, are provided in the Appendix C.
+Post-hoc baseline.
+SoftCAM was compared with widely used post-hoc methods (Appendix D), including CAM gradient-based approaches such as GradCAM [selvaraju2017grad] and LayerCAM [jiang2021layercam]; the CAM gradient-free method ScoreCAM [wang2020score]; and backpropagation-based techniques such as Integrated Gradients (Itgd Grad) [sundararajan2017axiomatic] and Guided Backpropagation (Guided BP) [springenberg2014striving].
+Each method was evaluated on its respective black-box CNN models.
 Evaluation metrics.
-Models were evaluated on both classification performance and explainability.
-For binary tasks, performance was measured using accuracy and AUC, while for multi-class tasks, accuracy and the quadratic Cohen’s kappa score were used.
-AUC reflects class separability, whereas the kappa score captures agreement beyond chance.
-To assess explainability, we employed several quantitative metrics: Top-k localization precision [18], activation precision [9, 43], activation consistency [18], and faithfulness [16, 39].
-We further extended activation precision to define activation sensitivity.
-For full descriptions of the explainability metrics, see Appendix B.
+The models were evaluated for predictive performance and explainability.
+For explainability, we used several quantitative metrics, including top-k localization precision [donteu2023sparse]; activation precision [barnett2021case], faithfulness [yeh2019fidelity], and activation consistency [donteu2023sparse], and we further extended activation precision to define activation sensitivity.
+Together, these metrics quantify both alignment with expert clinician annotations and alignment with the model’s actual decision-making process.
+Full metric descriptions are provided in Appendix E.
 4 Results
-5.1 Making black box CNNs self-explainable maintains classification performance
-Table 1: Classification performance for binary disease detection on the test sets. We denote the SoftCAM versions of ResNet and VGG with a .
+4.1 Making black box CNNs explainable maintains classification performance
+Table 1: Classification performance for binary disease detection on the test sets. SoftCAM variants of both CNNs are denoted by SC, with $\ell_{\lambda}$ indicating the applied penalty.
 Kaggle Fundus OCT retinal RSNA CXR
-Binary Multi-class Binary Multi-class Binary
-Acc. AUC Acc. $\kappa$ Acc. AUC Acc. $\kappa$ Acc. AUC
-VGG-16 $0.907$ $0.938$ $0.863$ $0.835$ $0.994$ $1.0$ $0.967$ $0.955$ $0.952$ $0.989$
-dense VGG $0.915$ $0.942$ $0.861$ $0.834$ $0.994$ $1.0$ $0.963$ $0.947$ $0.957$ $0.999$
-sparse VGG $0.911$ $0.938$ $0.859$ $0.827$ $0.988$ $0.999$ $0.947$ $0.929$ $0.953$ $0.990$
-ResNet-50 $0.899$ $0.923$ $0.850$ $0.800$ $0.994$ $0.999$ $0.970$ $0.963$ $0.953$ $0.988$
-dense ResNet $0.899$ $0.926$ $0.851$ $0.811$ $0.994$ $1.0$ $0.974$ $0.960$ $0.942$ $0.983$
-sparse ResNet $0.895$ $0.923$ $0.851$ $0.801$ $0.996$ $1.0$ $0.963$ $0.955$ $0.941$ $0.979$
-We first evaluated our method on clinically relevant classification tasks, including retinal disease classification from color fundus and OCT retinal images, as well as pneumonia detection from chest X-rays.
-For the fundus and OCT retinal datasets, both binary classification ({0} vs.
-{1-4} for fundus and Normal vs.
-Drusen for OCT) and multi-class classification tasks were considered, as reference labels were available.
-In contrast, the RSNA CXR dataset only included labels for pneumonia detection, restricting the evaluation to the binary task.
-For each CNN architecture, the “dense” model corresponds to our method without regularization ($\lambda_{1}=\lambda_{2}=0$), while the “sparse” model is obtained by applying a lasso penalty ($\lambda_{2}=0$) and choosing an appropriate value for $\lambda_{1}$ (e.g.
+Acc. AUC Acc. AUC Acc. AUC
+VGG-16 $0.907$ $0.938$ 0.994 1.0 $0.952$ $0.989$
+VGG-16SC $\mathbf{0.915}$ 0.942 0.994 1.0 0.957 0.999
+$\ell_{1}$-VGG-16SC $0.911$ $0.938$ $0.988$ $0.999$ $0.953$ $0.990$
+ResNet-50 0.899 $0.923$ $0.994$ $0.999$ 0.953 0.988
+ResNet-50SC 0.899 0.926 $0.994$ 1.0 $0.942$ $0.983$
+$\ell_{1}$-ResNet-50SC $0.895$ $0.923$ 0.996 1.0 $0.941$ $0.979$
+We first evaluated our method on clinically relevant binary classification tasks, including retinal disease classification from color fundus ({0} vs.
+{1-4}) and OCT retinal images (Normal vs.
+Drusen), as well as pneumonia detection from chest X-rays, using accuracy and AUC as primary metrics.
+For each CNN architecture, the $\text{}^{\text}{\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace SC\textquotedblright\ltxml@oqmark@close{}}$ model corresponds to our method without regularization ($\lambda_{1}=\lambda_{2}=0$), while its $\ell_{1}$ variant is obtained by applying a lasso penalty ($\lambda_{2}=0$) and choosing an appropriate value for $\lambda_{1}$ (e.g.
 $\lambda_{1}=1.10^{-6}$ for VGG and $\lambda_{1}=5.10^{-5}$ for ResNet on the fundus dataset).
-The sparsity parameter was selected by balancing classification accuracy and AUC on the validation set (see Appendix C).
-Our results show that SoftCAM models, which use explicit self-explainable class evidence maps, preserve classification performance comparable to their corresponding black-box counterparts (Tab.
-1).
-Moreover, introducing the Lasso regularization penalty on the class evidence map did not significantly degrade performance; in some cases, it even led to slight improvement.
+The sparsity parameter was selected by balancing classification accuracy and AUC on the validation set (Appendix F).
+Our results show that SoftCAM-based models (Tab.
+1), with explicit class evidence maps, preserve classification performance comparable to their corresponding black-box baselines.
+Moreover, introducing the lasso regularization on the class evidence map did not significantly degrade performance; in some cases, it even led to a slight improvement.
 These findings suggest that using convolutional layers in the classification head is an effective and promising approach for developing high-performing, self-explainable CNN models.
 4.2 SoftCAM provides inherently interpretable visual explanations
-Figure 2: Example explanations generated by different methods from ResNet-50.
-The first column shows disease images with reference annotations, indicated by green markers or bounding boxes.
-Each row, from top to bottom, corresponds to fundus, OCT, and Chest X-ray images, respectively.
-The next five columns present saliency maps generated by post-hoc explanation methods, gradient-free (ScoreCAM, LayerCAM) and gradient-based (GradCAM, Guided BP, Itgd Grad).
-The final two columns showcase our proposed inherently interpretable dense and sparse SoftCAM explanations.
-We qualitatively compared the evidence maps of SoftCAM variants with saliency maps generated by the five state-of-the-art CAM-based methods.
+We qualitatively compared the evidence maps of the SoftCAM variants with the saliency maps generated by the five state-of-the-art CAM and back-propagation-based methods.
 Overall, our method produced more visually interpretable maps with high evidence regions centered on annotated lesions (Fig.
 2).
-We observed that the regions highlighted by the sparse SoftCAM models are mostly a subset of those identified by the dense SoftCAM, reflecting the effect of the sparsity constraint.
-Additional results, including those for VGG-16 and other illustrative examples, are provided in Appendix D.1.
-On healthy images, sparse SoftCAM evidence maps exhibited overall more negative activations, in contrast to the positive activations observed on disease images.
-To assess this quantitatively, we computed the activation consistency [18], calculating the proportion of positive and negative activations across disease and healthy samples.
-These findings were consistent with the qualitative observations (e.g.
-dense vs.
-sparse SoftCAM on the fundus dataset using ResNet: $0.55$ vs.
-$0.27$ for the proportion of positive activation on disease images).
-For full analysis, see Appendix D.2.
+We observed that the regions highlighted by the sparse SoftCAM models are typically a subset of those identified by the base SoftCAM, reflecting the effect of the sparsity constraint in reducing irrelevant activations.
+Additional results, including those for VGG-16 and other examples, are provided in Appendix G.1.
+On healthy images, the sparse SoftCAM evidence maps exhibited overall more negative activations, in contrast to the positive activations observed on disease images.
+To assess this quantitatively, we computed the activation consistency score [donteu2023sparse], calculating the proportion of positive and negative activations in disease and healthy samples from the test set.
+These findings were consistent with qualitative visualization (e.g.
+sparse SoftCAM vs.
+SoftCAM on the fundus dataset using ResNet: $0.55$ vs.
+$0.27$ for the average proportion of positive activations on disease images).
+For a full analysis, see Appendix G.2.
+Figure 2: Example explanations generated by different methods from ResNet-50.
+The first column shows disease images with reference annotations (green markers or bounding boxes).
+Row from top to bottom corresponds to fundus, OCT, and Chest X-ray images, respectively.
+The next five columns present saliency maps from different post-hoc explanation methods.
+The last two columns show our proposed inherently interpretable dense and sparse SoftCAM explanations.
 4.3 SoftCAM provides localized and faithful explanations
-Figure 3: Quantitative evaluation of explanations generated by different methods.
-The first row shows the localization precision of the saliency maps on the Fundus and OCT datasets, evaluated against their respective ground truth.
-The second row presents the sensitivity analysis assessing the faithfulness of the generated explanations.
-Columns a,b show ResNet results, while c,d correspond to VGG.
-Higher precision means better localization; lower sensitivity implies more reliable explanations.
-[…]
-[…]
+Figure 3: Quantitative evaluation of explanations methods.
+The first row shows the localization precision of saliency maps on fundus and OCT datasets.
+The second row presents the sensitivity analysis assessing faithfulness.
+Columns a,b show ResNet results, and c,d correspond to VGG.
+Higher precision indicates better localization; lower sensitivity reflects more reliable explanations.
+To quantitatively assess the explanations provided by our SoftCAM-based evidence maps compared to post-hoc saliency methods, we first evaluated their localization precision, which measures how consistently the highlighted regions in the explanation maps align with clinician-annotated disease findings.
+Following [donteu2023sparse], we computed the Top-k (k=15) localization precision by upsampling each explanation map to the input resolution, splitting it into non-overlapping $33\times 33$ patches, and calculating the proportion of positively activated patches that overlap with ground truth annotations.
+Although inherently interpretable, SoftCAM-based explanations performed competitively overall in terms of localization precision (Fig.
+3).
+Notably, the sparse SoftCAM with the ResNet backbone outperformed all other methods with the highest top-k precision (Appendix G.3 and G.4), and ranked second only in top-3 precision on the fundus dataset (Fig.
+3a), behind Guided BP, which benefits from high-resolution saliency maps.
+Furthermore, the base and sparse SoftCAM typically achieved higher precision with fewer top-k regions, leading to fast convergence, particularly on the Fundus and OCT datasets.
+This suggests that sparse SoftCAM more consistently highlights fewer, yet truly relevant regions, whereas post-hoc methods produce broader and less specific activations, resulting in higher false-positive rates.
+Subsequently, we evaluated the faithfulness (also referred to as sensitivity) of the evidence maps generated by our SoftCAM-based explanations in comparison to post-hoc saliency maps.
+Sensitivity analysis evaluates how much the highly activated regions in an explanation map contribute to the model’s prediction [yeh2019fidelity], thereby assessing whether the highlighted areas actually influence the model’s decision-making process.
+To do this, we split the input images into non-overlapping $33\times 33$ patches, then progressively removed the top-ranked patches (based on attribution scores) and measured the relative change in model confidence.
+We conducted this evaluation on samples that were correctly predicted by both the black-box CNNs and their corresponding SoftCAM variants from the test sets.
+We found that the sparse SoftCAM generally outperformed other methods, notably on the OCT and RSNA datasets (Fig.
+3; Appendix G.3 and G.4).
+On the fundus dataset, both the base and sparse SoftCAM models performed slightly below the best-performing post-hoc methods, with Guided BP yielding the highest sensitivity scores, followed by Integrated Gradients (Fig.
+3).
+On the OCT dataset, the SoftCAM variants outperformed all post-hoc methods with the ResNet model, while with the VGG the sparse and base versions ranked second and third, respectively.
+On the RSNA dataset, the sparse SoftCAM achieved the highest sensitivity, surpassing all other methods, with the base SoftCAM ranking second with ResNet and third with VGG (Appendix G.4).
 4.4 Ridge regularization improves explanation for large disease regions
-Since the CXR dataset provided larger bounding boxes localizing disease regions, unlike the point-wise lesion annotations available in the fundus and OCT datasets, we computed activation precision [9, 43], which measures the proportion of the class-guided explanation that fall within the ground-truth bounding boxes, emphasizing precision by penalizing only false positives.
-However, it does not account for sensitivity or penalize false negatives.
-To address this limitation, we extended this metric to activation sensitivity (see Appendix B.2), which penalizes false negatives to better assess the explanation completeness, especially important in clinical imaging tasks where missing relevant regions can be critical, such as in multi-focal infectious diseases like pneumonia like pneumonia [37].
-We further investigated how different regularization strategies affect explanation quality.
-While Lasso regularization promoted sparsity by shrinking some activations to zeros, ridge regularization encouraged small (but nonzero) values, resulting in denser evidence maps.
-To evaluate this, we trained a ridge SoftCAM model ($\lambda_{1}=0$) and compared its performance to dense and sparse SoftCAM, as well as to the post-hoc explanation methods.
-The ridge penalty values were selected to balance classification performance ($\lambda_{2}=7.10^{-5}$ vs.
-$\lambda_{2}=2.10^{-4}$ for ResNet and VGG; see Appendix E.1).
-Under comparable classification performance (Acc.=$0.95$ for Ridge ResNet*, and VGG*), we found that all SoftCAM variants (dense, sparse, and ridge) generally outperformed the evaluated posthoc methods in both activation precision and activation sensitivity (Fig.
-4; Appendix E.2, E.3).
-Specifically, sparse SoftCAM achieved the highest activation precision, while ridge SoftCAM excelled in activation sensitivity.
-Dense SoftCAM consistently performed in between, underscoring the importance of balancing lasso and ridge regularization via ElasticNet to adapt to varying interpretability needs.
 Figure 4: Example of localization evaluation on the CXR dataset for pneumonia detection.
@@
 Ground-truth bounding boxes are overlaid on each map, with the top-right value indicating the activation precision, while the top-left value indicates the activation sensitivity.
-5.5 SoftCAM provides resource-efficient and faithful explanations for multi-class tasks
-Finally, we extended our method to the multi-class setting for retinal disease diagnosis.
-We retrained the same training setup as for the binary tasks, adjusting the output classes in the evidence layer to $5$ for DR grading (fundus dataset) and 4 for retinal disease classification (OCT dataset).
-Given the small size of retinal lesions, we used Lasso regularization, selecting $\lambda_{1}$ values that balanced performance (e.g.
-$\lambda_{1}=9.10^{-4}$ vs.
-$\lambda_{1}=3.10^{-6}$ for ResNet and VGG on the OCT dataset; Appendix F.1).
-Both dense and sparse models achieved performance comparable to their respective black-box baselines (Tab.1), with a slight improvement in Kappa on the fundus dataset when using the ResNet backbone.
-As no ground-truth lesion annotations were available for the multi-class tasks, we evaluated the faithfulness of the explanations by measuring their contribution to model predictions.
-For correctly classified test samples, we progressively removed top-k ($k=30$) ranked patches (based on the explanation maps; see Sec.5.3) and tracked the average drop in class confidence.
-In both tasks, the dense and sparse SoftCAM achieved superior performance, with sparse SoftCAM yielding the lowest area under the deletion curve, indicating the highest faithfulness (see Appendix F.2, F.3).
-Notably, the sparse SoftCAM produced class-wise explanations that aligned well with class model confidence, showing minimal evidence in healthy classes (Fig.5; Appendix F.4, F.5 for VGG and more examples).
-In the case of DR detection, a progressive disease, it is expected that images labeled with grade $x$, where $51<x<$ 5, may still exhibit features from earlier stages, consistent with explanations.
-Unlike post-hoc CAM-based methods, which require backpropagation or perturbation for each class, SoftCAM generates class-specific explanations during prediction in a single forward pass, making it more resource-efficient.
-Figure 5: Examples of multi-class explanations using ResNet.
-For a severe DR example from the Kaggle dataset, the first row shows class-specific dense SoftCAM evidence map explanations, while the second presents explanations from the sparse SoftCAM.
+Since the CXR dataset provided larger bounding boxes localizing disease regions, unlike the point-wise lesion annotations available in the fundus and OCT datasets, we computed activation precision [barnett2021case], which measures the proportion of the class-guided explanation that falls within the ground-truth bounding boxes, emphasizing precision by penalizing only false positives.
+However, it does not account for sensitivity by not penalizing false negatives.
+To address this, we extended this metric to activation sensitivity (Appendix E.3), which penalizes false negatives to better assess the explanation completeness, especially important in clinical imaging tasks where missing relevant regions can be critical.
+We further investigated how different regularization methods qualitatively and quantitatively affect explanations.
+While lasso regularization promoted sparsity by shrinking most false positive activations to zeros, ridge regularization encouraged small (but nonzero) values, resulting in denser evidence maps.
+To evaluate this, we trained a ridge SoftCAM model ($\lambda_{1}=0$) and compared its performance to the sparse SoftCAM, as well as to the post-hoc explanation methods.
+The ridge regularization strength was selected to balance predictive performance ($\lambda_{2}=7.10^{-5}$ vs.
+$\lambda_{2}=2.10^{-4}$ for ResNet and VGG; Appendix H.1).
+[…]
+4.5 SoftCAM provides faithful explanations for multi-class tasks
+We extended our method to multi-class settings, training the SoftCAM models from ResNet and VGG for DR detection (5 classes, Kaggle dataset) and retinal disease classification (4 classes, OCT dataset).
+The training setup remained consistent with the binary task, adjusting only the number of classes in the evidence layer and selecting appropriate lasso penalties.
+Detailed classification results, qualitative visualizations, and faithfulness analyses of the multi-class SoftCAM-based variants and their baseline are provided in the Appendix I.
+Briefly, both unregularized and sparse models achieved test accuracies comparable to their baseline (Acc $\approx 0.85$ vs.
+$0.97$ on Fundus and OCT), with SoftCAM, particularly the sparse variant, providing the most interpretable class-wise explanations with the best sensitivity.
 5 Discussion
-Here, we introduced SoftCAM, a straightforward yet effective approach for transforming black-box CNN models into inherently interpretable architectures.
-We tested SoftCAM on a diverse range of medical imaging tasks, including color fundus photographs, retinal OCT scans, and Chest X-rays for disease diagnosis.
-Importantly, SoftCAM-variants maintained performance comparable to that of the original CNN models for classification and generally outperformed post-hoc explainability techniques.
-SoftCAM produces explicit class evidence maps that directly contribute to the model’s prediction.
-This integration enables single forward-pass generation of explanations aligned with the classification output, resulting in resource-efficient and self-explainable CNNs.
-We evaluated our method for two widely used CNN backbones: ResNet-50 and VGG-16, assessing both classification performance and explainability.
-Despite some differences, both ResNet and VGG models employ large receptive fields, resulting in low-resolution feature maps.
-Consequently, the class-evidence layer operates on coarse feature maps, producing coarse-grained explanations.
-In the future, we could explore the integration of SoftCAM with other standard architectures like ViT [19].
-Our work presents a major step forward in the development of powerful self-explainable models, demonstrating that interpretable-by-design architectures can preserve, and in some cases even improve upon, the classification performance of state-of-the-art models by modifying standard, well-tested CNN architectures without the need for complicated additional concepts such as prototypes [15].
-Beyond performance, SoftCAM provides deeper insights into the model-decision-making process, offering a powerful tool for understanding mistakes and detecting spurious correlations, without relying on widely used post-hoc explanation methods.
-By leveraging ElasticNet regularization, which is task-specific, user can flexibly balance localization precision and sensitivity according to their application needs.
-This is especially relevant for CNN-based classifiers deployed in high-stakes decision-making contexts.
-We hope this contribution will pave the way toward designing more accurate and interpretable CNN models, ultimately fostering trust, adoption, and integration in critical real-world settings such as in medicine.
-Acknowledgments and Disclosure of Funding
-This project was supported by the Hertie Foundation, the German Science Foundation (Excellence Cluster EXC 2064 “Machine Learning—New Perspectives for Science”, project number 390727645).
-The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting KD.
-Appendix A Implementation Details
-A.1 Datasets
-We evaluate our approach on three publicly available medical imaging datasets spanning three different modalities: the Kaggle Diabetic Retinopathy (DR) [20], the Retinal OCT dataset [31], and the RSNA Chest X-ray (CXR) dataset [1].
+Here, we introduced SoftCAM, a lightweight architectural modification that makes standard black-box CNNs inherently interpretable without relying on post-hoc explanations.
+SoftCAM replaces the final pooling and FCL with 1×1 convolutions, producing class-specific evidence maps that directly inform predictions.
+This design also supports ElastiNet regularization on the evidence maps to improve explanations.
+As a result, SoftCAM produces aligned predictions and explanations in a single forward-pass, yielding resource-efficient, self-explainable CNNs.
+SoftCAM-based variants were validated on ResNet-50 and VGG-16 across three medical imaging modalities—fundus photographs, retinal OCT scans, and Chest X-rays—achieving classification performance comparable to their black-box baselines.
+For explainability, SoftCAM was benchmarked against five state-of-the-art post-hoc saliency methods—ScoreCAM, LayerCAM, GradCAM, Guided BP, and Itgd Grad—using qualitative and quantitative metrics that capture both alignment with human knowledge and alignment with the model’s true decision-making, demonstrating superior interpretability without compromising predictive performance.
+In particular, localization precision and activation sensitivity assess how well explanations match clinically relevant biomarkers using expert annotations, while faithfulness measures how highlighted regions truly influence the model’s predictions.
+Interestingly, our results revealed a discrepancy between human-aligned and model-aligned metrics (Fig.
+3): explanations with the best faithfulness did not always yield the strongest alignment with expert annotations.
+This suggests that models may rely on clinically hidden signals to humans and highlights the need for new evaluation metrics that jointly account for human knowledge and models’ internal reasoning.
+Despite its promising results, SoftCAM has some limitations that warrant further investigation.
+First, SoftCAM relies on the final low-resolution feature maps (e.g., 16×16 for standard CNNs), limiting the spatial granularity needed for fine-grained medical tasks that require lesion- or pixel-level precision.
+In our case, both ResNet and VGG models use large receptive fields, producing low-resolution feature maps.
+Consequently, the class-evidence layer generates coarse-grained explanations, though ElasticNet constraints help improve localization.
+Future work could focus on generating higher-resolution explanations with softCAM to overcome coarseness from downsampling in deep CNNs; integrating SoftCAM with other architectures like ViTs [dosovitskiy2020image]; extending it to tasks such as weakly supervised segmentations or object detection; and evaluating its utility in other high-stake domains, including agriculture or medical modalities such as dermoscopy for skin cancer detection or MRI for brain tumor detection.
+Finally, SoftCAM was compared only against a subset of relevant post-hoc methods, consistent with our primary goal of avoiding post-hoc explanations for CNNs classifiers.
+Future work could extend this evaluation to include additional post-hoc techniques and self-explainable models, such as part-prototype networks [chen2019looks], concept-based models [koh2020concept], or attention-based explainable architectures [djoumessi2025hybrid].
+[…]
+This project was supported by the Hertie Foundation, the German Science Foundation (Excellence Cluster EXC 2064 “Machine Learning—New Perspectives for Science”, project number 390727645; BE 5601/14-1, project number 571331899).
+The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Kerol Djoumessi.
+Appendix A Related work
+Explainable AI (XAI) methods for image analysis can be categorized into attribution and non-attribution-based approaches [he2022survey, hossain2023explainable].
+Attribution-based methods aim to explain “where” important features are located by generating saliency maps that assign importance scores to pixels or regions.
+In contrast, non-attribution methods focus on explaining “why” a decision was made, typically providing global explanations using approaches such as prototype- or concept-based explanations [chen2019looks, koh2020concept].
+Attribution-based methods usually provide post-hoc local explanations for black-box models after training.
+In contrast, non-attribution-based methods are generally inherently interpretable by design [chen2019looks, brendel2018bagnets, koh2020concept, bohle2022b], embedding transparency within their architecture.
+However, inherently interpretable models are typically model-agnostic and often face reduced performance, reflecting a persistent trade-off between interpretability and predictive performance [rudin2019stop].
+To bridge this gap, we propose SoftCAM, a protocol that makes traditional black-box CNNs inherently interpretable.
+Our approach builds on and generalizes prior work.
+aubreville2019transferability proposed a dual-branch architecture in which the first branch performed classification using a black-box CNN, while the second branch generated post-hoc explanations, requiring two separate forward passes.
+In the explanation branch, the global average pooling (GAP) layer is removed and the linear classifier is replaced with convolutional layers that share weights with the main classifier during inference, producing class-specific activation maps.
+In contrast, SoftCAM integrates both prediction and explanation within a single, end-to-end forward pass, eliminating the need for redundant computation and weights sharing.
+Similarly, donteu2023sparse leveraged a convolutional classifier to generate explicit class-evidence maps and applied a sparsity constraint to enhance interpretability in a self-explainable Bag-of-Local-Features model (BagNet) brendel2018bagnets.
+SoftCAM extends and generalizes this concept to standard black-box CNNs, broadening its applicability to various medical imaging tasks and modalities.
+Appendix B Datasets
+We evaluated our approach on three publicly available medical imaging datasets spanning three different modalities: the Kaggle Diabetic Retinopathy (DR) kaggle_dr_detection, the Retinal OCT dataset kermany2018identifying, and the RSNA Chest X-ray (CXR) dataset rsna_dataset.
   • Kaggle DR Dataset.
-This dataset comprises $88,702$ high-resolution retinal fundus images labeled for DR severity on a 5-point scale from 00 (No DR) to $4$ (Proliferative DR).
-After applying an automated quality filtering pipeline using an ensemble of EfficientNet models [55] trained on the ISBI2020222https://isbi.deepdr.org/challenge2.html challenge dataset, we retained $45,923$ images from $28,984$ subjects.
-The resulting class distribution was $73\%$ (class 0), $15\%$, $8\%$, $3\%$, and $1\%$.
+This dataset comprises $88,702$ high-resolution retinal fundus images labeled for DR severity on a 5-point scale from 0 (No DR) to $4$ (Proliferative DR).
+After applying an automated quality filtering pipeline using an ensemble of EfficientNet models trained on the ISBI2020111https://isbi.deepdr.org/challenge2.html challenge dataset, we retained $45,923$ images from $28,984$ subjects.
+The resulting class distribution was $73\%$, $15\%$, $8\%$, $3\%$, and $1\%$ for classes 0-4 respectively.
 For binary classification (early DR detection), we grouped class {0} vs.
@@
 $27\%$.
-Additionally, lesion annotations for $65$ images were obtained from [17] for evaluating the model’s explanations at localizing DR-related lesions.
+Additionally, lesion annotations for $65$ images were obtained from djoumessi2024inherently for evaluating the model’s explanations at localizing DR-related lesions.
   • Retinal OCT Dataset.
@@
 A separate test set of $1,000$ B-scans is provided.
-Following [16], we excluded low-resolution scans (width $\leq$ 496).
-As preliminary experiments showed that using the full dataset did not significantly improve performance, we subsampled the training set (by randomly removing half of the healthy images [16]) to $34,962$ scans ($8,616$ Drusen, $26,346$ Normal) for binary classification (Drusen vs.
+Following djoumessi2024actually, we excluded low-resolution scans (width $\leq$ 496).
+As preliminary experiments showed that using the full dataset did not significantly improve performance, we subsampled the training set (by randomly removing half of the healthy images following djoumessi2024actually) to $34,962$ scans ($8,616$ Drusen, $26,346$ Normal) for binary classification (Drusen vs.
 Normal), preserving the original class imbalance ($73\%$ vs $27\%$).
-Additionally, we used $40$ drusen-annotated B-scans from [16] to evaluate the model’s explanations at localizing drusen lesions.
+Additionally, we used $40$ drusen-annotated B-scans from djoumessi2024actually to evaluate the model’s explanations at localizing drusen lesions.
 For the multi-class classification task, the training was randomly reduced to $17,200$ images while maintaining the original class distribution: $45\%$ Normal, $34\%$ CNV, $10\%$ DME, and $9\%$ Drusen.
@@
 $40\%$ class distribution.
-Each dataset was split into training ($75\%$), validation ($10\%$), and test ($15\%$) sets, except for the Retinal OCT dataset, which followed an $80\%\text{-}20\%$ training-validation split, due to its predefined test set ( $250$ images per class).
-All training splits used in our experiments are provided in CSV format and publicly available via the project’s GitHub [^3] Code and CSV files are available at https://anonymous.4open.science/r/SoftCAM-E1A3/ repository.
-A.2 Baseline models
-The effectiveness of our method was evaluated using two widely adopted black-box CNN architectures: ResNet-50 [25] and VGG-16 [49].
+[…]
+Appendix C Implementation details
+C.1 Baseline models
+The effectiveness of our method was evaluated using two widely adopted black-box CNN architectures: ResNet-50 and VGG-16.
 These models were chosen due to their distinct architectures, such as depth, theoretical receptive field size, and classification head design, which allow for a broad assessment of our method’s generalizability.
-In both models, the standard classification head was replaced with our proposed convolutional evidence map layer to enable inherent interpretability.
-For ResNet50, we removed the global average pooling layer and final fully connected layer, substituting them with a class evidence layer consisting of $C$ convolutional filters ($1\times 1$, stride 1), where $C$ is the number of output classes.
+In both models, the standard classification head was systematically replaced with our proposed convolution-based evidence map layer to enable inherent interpretability.
+For ResNet50, we removed the global average pooling layer and final fully connected layer (FCL), substituting them with a class evidence layer consisting of $C$ convolutional filters ($1\times 1$, stride 1), where $C$ is the number of output classes.
 This layer directly produces class-specific evidence maps (Sec.
-3.2).
-For VGG-16, which uses a series of fully connected layers (FCLs) in its classifier head, each FCL was replaced by an equivalent $1\times 1$ convolutional layer.
-Specifically, an FCL of size $b_{1}\times b_{2}$ was transformed into a convolutional layer of size $b_{1}\times b_{2}\times 1\ times1$, preserving the original parameter count and model capacity.
-These architectural adjustments maintain model complexity and capacity while introducing interpretability directly into the classification mechanism.
-A.3 Data preprocessing
-Fundus images were preprocessed by cropping them to a square shape using a circle-fitting method as described in [38].
+2.2).
+For VGG-16, whose classifier head consist of several fully connected layers, each FCL was replaced by an equivalent $1\times 1$ convolutional layer.
+Specifically, an FCL of size $b_{1}\times b_{2}$ was transformed into a convolutional layer of size $b_{1}\times b_{2}\times 1\times 1$, preserving the original parameter count and model capacity.
+These architectural changes maintain model complexity and capacity while introducing interpretability directly into the classification mechanism.
+C.2 Data preprocessing and augmentation
+Fundus images were preprocessed by cropping them to a square shape using a circle-fitting method as described in Mueller_fundus_circle_cropping.
 All datasets were then resized to $512\times 512$ pixels, except for the retinal OCT dataset, which was resized to $496\times 496$ to better match its original lower resolution.
 Image intensities were normalized using the mean and standard deviation computed from the respective training sets.
-During training, consistent data augmentation strategies were applied across all datasets.
+During training, standard transformations were applied across all datasets.
 These included flipping, rotation, random cropping, and translation, each applied with a fixed probability.
 For the Kaggle dataset, which contains color fundus images, additional color augmentations were introduced to improve generalization.
-A.4 Training setup
-All models were obtained from Torchvision and initialized with pretrained ImageNet weights.
-They were subsequently fine-tuned on each dataset using a consistent training setup.
-Following [16, 18], we employed the cross-entropy loss function and optimized model parameters using stochastic gradient descent (SGD) with Nesterov momentum (momentum factor of $0.9$).
-The initial learning rate was set to $1.10\text{-}3$, and a clipped cosine annealing learning rate scheduling was applied with the minimum learning rate set to $1.10\text{-}4$.
-Weight decay was set to $5.10\text{-}4$.
-Training was conducted for 70 epochs with a mini-batch size of 16 on an NVIDIA A40 GPU using PyTorch [42].
-A.5 Baseline CAM-based methods
-Gradient-based methods primarily differ in how gradients are aggregated to compute importance weights, while gradient-free methods mainly vary in how the weights are computed.
+C.3 Training setup
+[…]
+Appendix D Baseline post-hoc methods
+SoftCAM was benchmarked against widely used post-hoc methods, spanning CAM-based and backpropagation-based approaches.
+CAM-based methods produce coarse-grained explanations due to their reliance on low-resolution convolutional feature maps, whereas backpropagation-based techniques provide fine-grained pixel-level attributions that perverse full spatial resolutions.
+CAM-based methods include gradient-based approaches such as GradCAM and LayerCAM, as well as the gradient-free method ScoreCAM.
+Backpropagation-based techniques include Integrated Gradients and Guided Backpropagation.
+Gradient-based methods primarily differ in how they aggregate gradients to compute importance weights, while gradient-free methods vary in how these weights are estimated without backpropagation.
+Guided Backpropagation and Integrated Gradients have consistently performed well in generating saliency maps to explain black-box CNN classifiers on retinal images ayhan2022clinical; djoumessi2024inherently, while GradCAM has shown strong localization performance for chest X-ray interpretation saporta2022benchmarking.
+Below is a brief description of the five post-hoc methods used.
+CAM
 ScoreCAM
-[56].
+wang2020score.
 A gradient-free method that eliminates the need for gradient information by assessing the importance of each activation map based on its forward-pass contribution to the target class score, and produces the final output via a weighted sum of these maps.
 LayerCAM
-[30].
+jiang2021layercam.
 A gradient-based method that generates class activation maps by leveraging the element-wise product of ReLU-activated gradients and feature maps at any convolutional layer, enabling fine-grained, spatially precise visual explanations without requiring global average pooling.
 GradCAM
-[47].
+selvaraju2017grad.
 A gradient-based approach that uses the gradients of the target class flowing into the final convolutional layer to produce a coarse localization map, highlighting important regions in the image by upsampling the resulting map.
 Guided backpropagation (Guided BP)
-[51].
+springenberg2014striving.
 A gradient-based approach that modifies the standard backpropagation process to propagate only positive gradients through positive activations, producing fine-grained visualizations that highlight features strongly activating specific neurons in relation to the target output.
 Integrated Gradient (Itgt Grad.)
-[53].
+sundararajan2017axiomatic.
 A gradient-based method that attributes model predictions to input features by computing the path integral of gradients along a straight-line path from a baseline to the actual input, yielding fine-grained explanations.
-Appendix B Explainability metrics
-B.1 Activation consistency
-The activation consistency [18] quantifies how well local explanations (e.g., individual activations within explanation maps) globally reflect the disease and healthy samples across a dataset.
-Specifically, it measures whether the activation patterns in the explanation maps consistently reflect the underlying disease or healthy class labels.
-Following [18], we evaluated activation consistency by computing the proportion of positive activations (indicative of disease evidence) in saliency maps of disease samples, and negative activations (indicative of the absence of disease) in those of healthy samples.
-These proportions were calculated over the test set to assess whether the heatmaps consistently highlight pathological features in diseased cases and suppress activations in healthy ones.
-This metric thus captures the alignment between the semantic meaning of activations and the ground truth labels, offering a dataset-level evaluation of the coherence of local explanations with the global classification objective.
-B.2 Activation precision and activation sensitivity
-Let $\mathcal{X}=\{\bf{X}\}_{i=1}^{n}$ denote a set of input images, =$\mathcal{M}=\{\bf{M}\}_{i=1}^{n}$ the corresponding binary segmentation masks, and $\mathcal{S}=\{\bf{S}\}_{i=1}^{n}$ the associated explanation or saliency maps generated by any method.
-Activation precision measures the proportion of the saliency map’s positive mass that lies within the annotated region (the segmentation mask) [9, 43].
-To compute it, saliency maps are first preprocessed by setting negative values to zero while retaining all positive values.
-This highlights how much of the explanation signal aligns with human-annotated ground truth, effectively quantifying the precision of an explainability method.
-The activation precision is defined as:
-$AP(\mathcal{M},\mathcal{S})=\frac{\sum_{M,S}\sum_{i,j}M_{i,j}.S_{i,j}}{\sum_{i,j}S_{i,j}}.$ (6)
-However, activation precision does not penalize false-negative (i.e.
+ScoreCAM and LayerCAM were implemented with TorchCAM torcham2020, while the other methods were implemented from Captum kokhlikyan2020captum.
+Appendix E Explainability metrics
+For binary tasks, performance was evaluated using accuracy and AUC, while for multi-class tasks, accuracy and the quadratic Cohen’s kappa score ($\kappa$) were used.
+The AUC measures class separability, whereas the kappa score captures the agreement beyond chance.
+Explainability was assessed using several quantitative metrics, including activation consistency donteu2023sparse, top-k localization precision donteu2023sparse, activation precision barnett2021case, further extended to activation sensitivity, and faithfulness yeh2019fidelity.
+Together, these metrics quantify both alignment with expert clinical knowledge (top-k localization, activation precision, and activation sensitivity) and alignment with the model’s true decision-making process (faithfulness).
+E.1 Activation consistency
+The activation consistency donteu2023sparse measures how well local explanations (i.e., positive or negative activation in attribution maps) reflect the true disease or healthy labels across a dataset.
+Intuitively, an interpretable model should consistently show positive activation regions associated with pathology for disease samples, while producing minimal or negative activations for healthy samples.
+This metric therefore assesses whether explanation maps globally align with the semantic meaning implied by ground-truth labels.
+Following donteu2023sparse, activation consistency is computed as the proportion of positive activations in the attribution maps of disease samples and the proportion of negative activations in the attribution maps of healthy samples, averaged over the test set.
+A higher score indicates more coherent and label-aligned explanations.
+[…]
+[…]
+corresponding to the proportion of positive and negative activations in the attribution map. Activation consistency is then defined as
+$AC_{+}=$ $\frac{1}{|\mathcal{D}_{\text{disease}}|}\sum_{i\in\mathcal{D}_{\text{disease}}}\mathbbm{1}_{+}(A_{i});$ (5)
+$AC_{-}=$ $\frac{1}{|\mathcal{D}_{\text{healthy}}|}\sum_{i\in\mathcal{D}_{\text{healthy}}}\mathbbm{1}_{-}(A_{i}),$ (6)
+where $AC_{+}$ measures how consistently explanations highlight pathological regions in disease samples, $AC_{-}$ measures how consistently they suppress activations in healthy samples.
+This provides a dataset-level assessment of whether local explanations globally support the model’s classification behavior.
+E.2 Top-k localization precison
+Top-k localization precision donteu2023sparse measures how well an explanation map highlights clinically relevant regions that overlap with ground-truth annotations.
+Specifically, it computes the proportion of the top-k positively activated regions that coincide with annotated areas obtained from clinicians.
+Given an explanation map, we first upsampled it to the original input resolution and split into non-overlapping patches of size $33\times 33$.
+Each patches is assigned a saliency score equal to the average activation within that patch.
+The top-k ($k\in[1,30]$) most salient patches are then selected, and the metric computes the fraction of these patches that overlap with annotated ground-truth regions.
+This generalizes the “pointing game” metric zhang2018top, which only considers the single most salient region (top-1), making it better suited for medical imaging tasks where disease-relevant features (e.g., retinal lesions or other pathological markers) are often spatially distributed across the image.
+Formally, let $S$ denote the upsampled explanation map and let the image be partitioned into $P$ non-overlapping patches $\{P_{1},\ldots,P_{p}\}$. The average activation of each patch is
+$a_{p}=\frac{1}{|P_{p}|}\sum_{(x,y)\in P_{p}}S(x,y).$
+[…]
+[…]
+The Top-k localization precision is then defined as
+[…]
+A higher score indicates that the explanation consistently highlights clinically meaningful regions identified by experts.
+E.3 Activation precision and activation sensitivity
+Let $\mathcal{X}=\{\bf{X}\}_{i=1}^{n}$ denote a set of input images, $\mathcal{M}=\{\bf{M}\}_{i=1}^{n}$ their corresponding binary segmentation masks, and $\mathcal{S}=\{\bf{S}\}_{i=1}^{n}$ the explanation or saliency maps produced by a given method.
+Activation precision measures how much of the positive evidence highlighted by an explanation map falls within the annotated ground-truth region barnett2021case.
+Before computing the metric, saliency maps are preprocessed by thresholding negative values to zero, $\mathbf{S}_{i}^{+}=\max(\mathbf{S}_{i},0)$, ensuring that only positive evidence contributes to the evaluation.
+For each sample, activation precision (AP) is defined as the proportion of positive activation mass contained inside the annotation mask:
+$\mathbf{AP}_{i}=\frac{\sum_{p}\mathbf{S}_{i}^{+}(p)\textbf{M}_{i}(p)}{\sum_{p}\mathbf{S}_{i}^{+}(p)+\epsilon},$
+where $p$ indexes spatial locations and $\epsilon$ is a small constant preventing division by zero. The dataset-level activation precision is then obtained by averaging over all samples:
+$\textbf{AP}=\frac{1}{n}\sum_{i=1}^{n}\textbf{AP}_{i}.$ (8)
+This metric captures how precisely the explanations signal aligns with expert annotation, providing a clinically meaningful measure of the quality of an explanation method.
+However, activation precision does not penalize false negatives (i.e.
 missed relevant regions).
-To address this, we introduce activation sensitivity, which captures the completeness of the explanation by evaluating how much of the annotated region is covered by the saliency map.
-The activation precision is defined as:
-$AS(\mathcal{M},\mathcal{S})=\frac{\sum_{M,S}\sum_{i,j}M_{i,j}.S_{i,j}}{\sum_{i,j}M_{i,j}}.$ (7)
-Unlike activation precision, activation sensitivity penalizes low saliency values within the mask.
-For example, if $M_{i,j}=1$ but $10<S_{i,j}<$ 1, the low activation will contribute little to the numerator, reflecting reduced confidence in that region.
-This makes activation sensitivity especially relevant in clinical tasks where completeness is critical, such as identifying multi-focal infectious diseases like pneumonia [37].
-B.3 Top-k localization precison
-Top-k localization precision [18] measures the ability of an explanation map to correctly highlight salient regions that overlap with ground-truth annotations.
-Specifically, it quantifies the proportion of the top-k positively activated regions within an explanation that match with annotated areas.
-In our implementation, each explanation map is first upsampled to the input resolution and then split into non-overlapping patches of size $33\times 3$ 33.
-These patches are ranked based on their average activation, and the top-k (k=30) most salient patches are selected.
-The precision is then computed as the fraction of these patches that overlap with the annotated ground-truth regions.
-This metric can be viewed as a generalization of the pointing game metric [60], where only the single most activated region (top-1) is considered, to multiple regions, making it more suitable for medical imaging tasks.
-In such contexts, disease-relevant features (e.g., retinal lesions or pathological markers) are often spatially distributed across the image, rather than confined to a single localized area.
-B.4 Failthfulness
-Faithfulness, also referred to as sensitivity or fidelity [39], is a widely used metric to evaluate how accurately an explanation reflects the model’s true decision-making process.
-It assesses whether the importance scores (attributions) assigned to input features correspond to the actual impact of those features on the model’s prediction.
-In our implementation, we focus on correctly classified samples from the test set.
-For each, the corresponding explanation map is upsampled to the input resolution and split into non-overlapping patches of size $33\times 3$ 33.
-These patches are ranked based on their mean activation values, and the top-k (k=30) most salient patches are iteratively occluded.
-After each occlusion step, we recorded the relative drop in the model’s confidence score for the predicted class.
-This process yields a deletion curve, from which we compute the Area Under the Deletion Curve (AUDC).
-A lower AUDC indicates a more faithful explanation, as it reflects a greater decline in model confidence when the most important regions (as indicated by the explanation map) are removed, suggesting that those regions were indeed critical to the model’s prediction.
-Appendix C Effect of Lasso regularization on model performance for the binary tasks
-The Lasso regularization coefficient $\lambda_{1}$ in Eq.
+To address this, we introduce activation sensitivity, which captures the completeness of the explanation by computing the fraction of annotated regions that are covered by the saliency map.
+We defined per-sample activation sensitivity (AS) by computing the proportion of mask pixels that receive any positive activation:
+$\mathbf{AS}_{i}=\frac{\sum_{p}\mathbf{S}_{i}^{+}(p)\textbf{M}_{i}(p)}{\sum_{p}\mathbf{M}_{i}(p)+\epsilon},$
+The dataset-level activation sensitivity is obtained by averaging over samples:
+$\textbf{AS}=\frac{1}{n}\sum_{i=1}^{n}\textbf{AS}_{i}.$ (9)
+Intuitively, $\textbf{AS}_{i}$ is the fraction of the annotated region that the explanation covers with higher values meaning fewer missed regions.
+Unlike activation precision, activation sensitivity penalizes weak activations inside the annotated region.
+For example, if $M_{i}(p)=1$ and $0<S_{i}(p)<1$, the low activation contributes only minimally to the numerator, indicating reduced confidence in that clinically relevant area.
+This makes activation sensitivity particularly valuable in settings where complete lesion coverage is essential.
+E.4 Failthfulness
+Faithfulness, also referred to as sensitivity or fidelity yeh2019fidelity, evaluates how well an explanation reflects the model’s true decision-making process.
+It assesses whether the importance assigned to input features corresponds to their actual influence on the model’s prediction.
+In our implementation, we evaluated faithfulness on correctly classified test samples.
+Each explanation map is upsampled to the input resolution and split into non-overlapping $33\times 33$ patches, which are ranked by mean activation values.
+The top-k most salient patches are iteratively occluded, and after each step we record the relative drop in the model’s confidence for the predicted class.
+This produces a deletion curve from which we computed the Area Under the Deletion Curve (AUDC).
+A lower AUDC indicates a more faithful explanation, as it reflects a sharper decrease in confidence when the regions deemed important are removed, showing that these regions effectively contributed to the model’s decision.
+Appendix F Sparsity regularization selection for the binary tasks
+The sparsity regularization coefficient $\lambda_{1}$ in Eq.
 4 controls the sparsity of the class evidence map, encouraging the model to localize disease regions with high precision.
-For each task, $\lambda_{1}$ was selected based on a trade-off between accuracy and AUC on the corresponding validation set, choosing the highest values for which classification performance did not degrade significantly.
-Figure 6: Model selection on validation sets under varying Lasso regularization strengths.
+For each task, $\lambda_{1}$ was selected based on a trade-off between accuracy and AUC on the corresponding validation set, choosing the highest values for which classification performance did not degrade significantly (Fig.
+9).
+Figure 5: Model selection on validation sets under varying sparsity regularization strengths.
 The regularization coefficient $\lambda$ influences model performance, with notable effects on some datasets but minimal impact on the OCT dataset.
 The red markers indicate the selected $\lambda$ values, chosen to balance sparsity and classification performance.
-Appendix D Additionnal Results
-D.1 SoftCAM provides inherently interpretable visual explanations
-Figure 7: Example explanations generated by different methods from VGG-16.
+Appendix G Additionnal Results
+Visualizing explanations.
+The evidence map generated by SoftCAM is upsampled to the input resolution for visualization.
+Like most CAM-based methods, such as GradCAM, ScoreCAM, and LayerCAM that operate on the final convolutional layer, SoftCAM’s explanations are limited by the resolution of the backbone (e.g., 16×16 for VGG-16/ResNet-50 with 512×512 input) due to pooling and striding, leading to lower-resolution saliency maps.
+However, by introducing the class evidence and classification layer directly on top of the low-resolution features and applying regularization, SoftCAM improves spatial precision.
+In contrast, gradient-based methods like Integrated Gradients sundararajan2017axiomatic and Guided Backpropagation springenberg2014striving produce high-resolution saliency maps by computing pixel-level gradients, which may lead to noisy maps, especially when the region of interest spans a broader area, as commonly observed in Chest X-ray images.
+Comparison with other approaches.
+Unlike post-hoc attribution-based approaches, SoftCAM is inherently interpretable from the classification layer and maintains performance comparable to its black-box counterpart, without a significant trade-off, even when regularization is applied to enhance explainability.
+Compared to donteu2023sparse, SoftCAM extends from interpretable bag-of-local models to general black-box CNN architectures and generalizes the regularization from lasso to ElasticNet, with extensive evaluations across multiple datasets using a broad range of explanability metrics.
+Compared to aubreville2019transferability, our method is trained end-to-end and does not require post-hoc processing, weight sharing between branches, or an additional forward pass to generate explanations.
+G.1 SoftCAM provides inherently interpretable visual explanations
+Figure 6: Explanations generated by different methods from VGG-16.
 The first column shows disease images with reference annotations, indicated by green markers or bounding boxes.
 Each row, from top to bottom, corresponds to fundus, OCT, and Chest X-ray images, respectively.
-The next five columns present saliency maps generated by post-hoc explanation methods, gradient-free (ScoreCAM, LayerCAM) and gradient-based (GradCAM, Guided BP, Itgd Grad).
-The final two columns showcase our proposed inherently interpretable dense and sparse SoftCAM explanations.
-Figure 8: Additional example explanations of disease images from the ResNet model.
-Figure 9: Additional example explanations of disease images from the VGG model.
-D.2 Activation consistency
-We quantify activation consistency only for the SoftCAM variants, as post-hoc methods are not inherently explainable, meaning their explanations do not directly influence the model’s decision-making process.
+The next five columns present saliency maps generated by post-hoc explanation methods.
+The final two columns showcase our proposed inherently interpretable SoftCAM explanations.
+Additional qualitative examples
+Figure 7: Additional qualitative explanations from the ResNet model.
+Figure 8: Additional qualitative explanations from the VGG model.
+G.2 Activation consistency
+We quantified activation consistency only for the SoftCAM variants, as post-hoc methods are not inherently explainable, meaning their explanations do not directly influence the model’s decision-making process.
 The results align well with qualitative visualizations.
 On the Fundus dataset, the sparse SoftCAM model exhibits a higher proportion of positive activations with the ResNet backbone, attributed to reduced false positives from the dense model, and fewer negative activations, reflecting the suppression of low-importance activations to zero.
-On the VGG backbone, regularization primarily reduces false-positive activations from the dense model but leads to a slight increase in activations on healthy samples.
-Similar result can be observed on the RSNA dataset.
-On the OCT dataset, the dense SoftCAM with the ResNet backbone generally produces coarse-grained evidence around lesion areas.
+On the VGG backbone, regularization primarily reduces false-positive activations from the unregularized model but leads to a slight increase in activations on healthy samples.
+A similar result can be observed on the RSNA dataset.
+On the OCT dataset, the unregularized SoftCAM with the ResNet backbone generally produces coarse-grained evidence around lesion areas.
 In contrast, the sparse variant refines these explanations, resulting in lower positive and negative activations across both disease and healthy samples, suggesting more selective and focused localization.
-However, with the VGG backbone, a higher proportion of negative activations is observed, reflecting the impact of the regularization strength, highlighting the importance of appropriately tuning this parameter for different architectures.
+However, with the VGG backbone, a higher proportion of negative activations is observed, reflecting the impact of the regularization strength—highlighting the importance of appropriately tuning this parameter for different architectures.
 Table 2: Activation consistency on the ResNet model.
@@
 $\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$ $\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$ $\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$
-dense SoftCAM $0.28\pm 0.1$ $0.86\pm 0.1$ $0.30\pm 0.1$ $0.85\pm 0.1$ $0.75\pm 0.1$ $0.47\pm 0.1$
+SoftCAM $0.28\pm 0.1$ $0.86\pm 0.1$ $0.30\pm 0.1$ $0.85\pm 0.1$ $0.75\pm 0.1$ $0.47\pm 0.1$
 sparse SoftCAM $0.55\pm 0.2$ $0.76\pm 0.2$ $0.23\pm 0.1$ $0.83\pm 0.1$ $0.79\pm 0.1$ $0.45\pm 0.1$
@@
 $\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$ $\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$ $\mathbf{r_{LG}^{+}}\uparrow$ $\mathbf{r_{LG}^{-}}\uparrow$
-dense SoftCAM $0.32\pm 0.2$ $0.93\pm 0.1$ $0.75\pm 0.11$ $0.51\pm 0.1$ $0.75\pm 0.1$ $0.51\pm 0.1$
+SoftCAM $0.32\pm 0.2$ $0.93\pm 0.1$ $0.75\pm 0.11$ $0.51\pm 0.1$ $0.75\pm 0.1$ $0.51\pm 0.1$
 sparse SoftCAM $0.28\pm 0.2$ $0.94\pm 0.1$ $0.35\pm 0.14$ $0.95\pm 0.1$ $0.35\pm 0.1$ $0.95\pm 0.1$
@@
 Nevertheless, the activation consistency metric aligns well with the qualitative explanations, generally capturing the impact of regularization across the dataset for a given architecture.
-D.3 Precision and sensitivity analysis
+G.3 Precision and sensitivity analysis
 We quantitatively evaluate the explanations generated by various methods using the ResNet and VGG backbones on the RSNA dataset.
-With the ResNet model, the dense SoftCAM achieves the highest localization precision, whereas the sparse SoftCAM yields the best results in terms of sensitivity.
+With the ResNet model, the unregularized SoftCAM achieves the highest localization precision, whereas the sparse SoftCAM yields the best results in terms of sensitivity.
 This discrepancy underscores the importance of developing evaluation metrics that balance human-aligned localization quality with model fidelity, capturing both interpretability and decision relevance.
 Figure 9: Precision vs. sensitivity analysis on the RSNA dataset. Quantitative evaluation of explanations generated by different methods from the ResNet and VGG models on the RSNA dataset.
-D.4 SoftCAM provides localized and faithful explanations
-Table 4: Top-k localization precision and sensitivity.
+G.4 SoftCAM provides localized and faithful explanations
+Alongside the visualization comparing SoftCAM variants with post-hoc methods, we provide the corresponding quantitative metrics, reporting both top-k precision and faithfulness.
+Table 4: Top-15 localization precision and sensitivity.
+Sensitivity is quantified as the Area Under the Deleted Curve (AUDC), where lower values indicate greater faithfulness.
+For precision, higher values indicate better alignment between saliency maps and ground truth annotations.
+We refer to AUDC as “Del” and Top-K as “Top”.
+ResNet (Top $\uparrow$, Del $\downarrow$) VGG (Top $\uparrow$, Del $\downarrow$)
+Fundus OCT RSNA Fundus OCT RSNA
+Top Del Top Del Top Del Top Del Top Del Top Del
+ScoreCAM $0.22$ $0.77$ $0.12$ $0.84$ $0.72$ $0.98$ $0.30$ $0.78$ $0.12$ $0.76$ $\bf{0.74}$ $0.93$
+LayerCAM $0.25$ $0.76$ $0.11$ $0.85$ $0.75$ $0.97$ $0.30$ $0.76$ $0.11$ $0.76$ $0.72$ $0.91$
+GradCAM $0.37$ $0.75$ $0.15$ $0.84$ $0.75$ $0.97$ $\bf{0.65}$ $0.79$ $0.15$ $0.77$ $0.72$ $0.92$
+Guided BP $0.38$ $\bf{0.69}$ $0.36$ $0.80$ $0.60$ $0.98$ $0.48$ $\bf{0.72}$ $0.36$ $\bf{0.65}$ $0.66$ $0.92$
+Itgd Grad $0.34$ $0.73$ $0.30$ $0.82$ $0.56$ $0.98$ $0.40$ $0.75$ $0.30$ $0.77$ $0.61$ $0.93$
+SoftCAM $0.40$ $0.79$ $0.46$ $0.84$ $0.74$ $\bf{0.95}$ $0.54$ $0.73$ $0.72$ $0.68$ $\bf{0.74}$ $0.92$
+sparse SoftCAM $\bf{0.52}$ $0.78$ $\bf{0.86}$ $\bf{0.57}$ $\bf{0.78}$ $\bf{0.95}$ $\bf{0.65}$ $0.78$ $\bf{0.82}$ $0.66$ $0.71$ $\bf{0.90}$
+Table 5: Top-k localization precision and sensitivity, $k=30$.
 Sensitivity is quantified as the Area Under the Deleted Curve (AUDC), where lower values indicate greater faithfulness—that is, a larger drop in the model’s confidence when the most relevant patches are removed.
 For precision, higher values indicate better alignment between saliency maps and ground truth annotations.
-We refer to AUDC as “Del” and Top-K as “Top” with $K=30$.
-ResNet (Topk $\uparrow$, AUDC $\downarrow$) VGG (Topk $\uparrow$, AUDC $\downarrow$)
+ResNet (Top $\uparrow$, Del $\downarrow$) VGG (Top $\uparrow$, Del $\downarrow$)
 Fundus OCT RSNA Fundus OCT RSNA
@@
 LayerCAM $0.22$ $0.65$ $0.08$ $0.74$ $0.65$ $0.97$ $0.23$ $0.64$ $0.08$ $0.56$ $\bf{0.65}$ $0.84$
-GradCAM $0.37$ $0.64$ $0.14$ $0.73$ $0.67$ $0.95$ $0.65$ $0.68$ $0.14$ $0.58$ $0.61$ $0.86$
+GradCAM $0.37$ $0.64$ $0.14$ $0.73$ $0.67$ $0.95$ $\bf{0.65}$ $0.68$ $0.14$ $0.58$ $0.61$ $0.86$
 Guided BP $0.30$ $\bf{0.57}$ $0.23$ $0.68$ $0.55$ $0.97$ $0.43$ $0.57$ $0.23$ $\bf{0.40}$ $0.58$ $0.85$
-Itgd Grad. $0.28$ $0.63$ $0.20$ $0.70$ $0.52$ $0.98$ $0.33$ $\bf{0.62}$ $0.2$ $0.51$ $0.55$ $0.88$
-dense SoftCAM $0.39$ $0.69$ $0.46$ $0.61$ $0.65$ $\bf{0.92}$ $0.54$ $0.63$ $0.72$ $0.45$ $0.64$ $0.84$
+Itgd Grad $0.28$ $0.63$ $0.20$ $0.70$ $0.52$ $0.98$ $0.33$ $\bf{0.62}$ $0.2$ $0.51$ $0.55$ $0.88$
+SoftCAM $0.39$ $0.69$ $0.46$ $0.61$ $0.65$ $\bf{0.92}$ $0.54$ $0.63$ $0.72$ $0.45$ $0.64$ $0.84$
 sparse SoftCAM $\bf{0.52}$ $0.68$ $\bf{0.86}$ $\bf{0.31}$ $\bf{0.73}$ $0.93$ $\bf{0.65}$ $0.674$ $\bf{0.82}$ $0.43$ $0.63$ $\bf{0.82}$
-Appendix E Activation precision and sensitivity on the RSNA dataset
-E.1 Lasso vs Ridge penalty
-Figure 11: Model selection on validation sets under varying Lasso and Ridge regularization strengths.
-The regularization coefficients $\lambda_{1}$ and $\lambda_{2}$ influence model performance.
+Appendix H Activation precision and sensitivity on the RSNA dataset
+H.1 Lasso vs Ridge penalty
+Figure 10: Model selection on validation sets under varying values.
+The regularization strengths $\lambda_{1}$ and $\lambda_{2}$ influence model performance.
 The red markers indicate the selected regularization values, chosen to balance classification performance.
-E.2 Activation precision vs. activation sensitivity
+H.2 Activation precision vs. activation sensitivity
 Table 6: Activation Precision (AP) vs.
 Activation Sensitivity (AS) for different SoftCAM variants and baseline post-hoc methods.
-The dense SoftCAM consistently lies between the lasso and ridge variants, highlighting the importance of balancing the two regularization terms to achieve an optimal trade-off between precision and completeness in the explanations.
+The base SoftCAM consistently lies between the lasso and ridge variants, highlighting the importance of balancing $\ell_{1}$ and $\ell_{2}$ values to achieve an optimal trade-off between precision and completeness.
 ResNet VGG
@@
 Itgd Grad. $0.286$ $0.040$ $0.322$ $0.039$
-dense SoftCAM $0.526$ $0.251$ $0.461$ $0.355$
+SoftCAM $0.526$ $0.251$ $0.461$ $0.355$
+ridge SoftCAM $0.440$ $\bf{0.316}$ $0.412$ $\bf{0.396}$
 sparse SoftCAM $\bf{0.654}$ $0.182$ $\bf{0.519}$ $0.320$
-lasso SoftCAM $0.440$ $\bf{0.316}$ $0.412$ $\bf{0.396}$
-E.3 More examples: activation precision vs activation sensitivity
-Figure 12: Additional examples of localization evaluation on the RSNA dataset for pneumonia detection.
-Each column shows explanation maps generated by different methods.
-Ground-truth bounding boxes are overlaid on each map, with the top-right value indicating the activation precision, while the top-left value indicates the activation sensitivity.
-The high precision from the lasso model and the more complete explanations from the ridge model emphasize the importance of balancing the two regularization terms to achieve an optimal trade-off.
-Appendix F Multi-class analysis
-F.1 Regularization
+H.3 More examples: activation precision vs activation sensitivity
+Figure 11: Localization precision and sensitivity for pneumonia detection.
+Each column shows explanations generated by different methods.
+Ground-truth bounding boxes are drawn on each map, with the top-right value indicating the activation precision, while the top-left value indicates the activation sensitivity.
+This shows the trade-off between ridge and lasso regularization.
+Appendix I Multi-class analysis
+We extended our method to the multi-class setting for retinal disease diagnosis, using the same training setup as in the binary tasks.
+The only modification was adjusting the number of output classes in the evidence layer to 5 for DR grading (fundus dataset) and 4 for retinal disease classification (OCT dataset, and selecting appropriate lasso penalties (e.g.
+$\lambda_{1}=9.10^{-4}$ vs.
+$\lambda_{1}=3.10^{-6}$ for ResNet and VGG on the OCT dataset).
+I.1 Regularization
 Given the small size of retinal lesions, we used Lasso regularization, selecting $\lambda_{1}$ values that balanced performance (Fig. 12)
-Figure 13: Model selection on validation sets under varying Lasso regularization strengths.
+Figure 12: Model selection on validation sets under varying sparsity values.
 The regularization coefficients $\lambda_{1}$ influence model performance.
 The red markers indicate the selected regularization values to balance classification performance.
-F.2 Faithfulness
-As no ground-truth lesion annotations were available for the multi-class tasks, we evaluated the faithfulness of the explanations by measuring their contribution to model predictions.
-For correctly classified test samples, we progressively removed top-k ($k=30$) ranked patches (based on the explanation maps) and tracked the average drop in class confidence (Fig.
-14).
-Figure 14: Sensitivity analysis.
-F.3 Area Under the Deleted Curve
-The area under the deletion curve (AUDC) was computed from the sensitivity analysis (Fig.
-14).
+I.2 Area Under the Deleted Curve
+The relative area under the deletion curve (AUDC) was computed from the sensitivity analysis by occluding the top-30 patches ranked by importance in the explanation map (Fig.
+LABEL:fig5:multi-class-sensitivity).
 In both tasks, the dense and sparse SoftCAM achieved superior performance, with sparse SoftCAM yielding the lowest AUDC, indicating the highest faithfulness (Tab.
-6).
+7).
 Table 7: Area Under the Deleted Curve (AUDC $\downarrow$).
@@
 Guided BP $0.899$ $0.793$ $0.905$ $0.823$
-Itgd Grad. $0.907$ $0.821$ $0.901$ $0.833$
+Itgd Grad $0.907$ $0.821$ $0.901$ $0.833$
 dense SoftCAM $0.870$ $0.825$ $0.905$ $0.826$
 sparse SoftCAM $\bf{0.856}$ $\bf{0.609}$ $0.882$ $\bf{0.806}$
-F.4 Qualitative explanation on retinal fundus images
-For the multi-class tasks on DR detection from fundus images, SoftCAM variants produced more focused and class-consistent explanations.
-In addition to the sparse and dense evidence maps, we also provide visualizations for post-hoc methods.
+I.3 Qualitative explanation on retinal fundus images
+For the multi-class DR detection tasks on fundus images, SoftCAM variants produced more focused and class-consistent explanations.
+Alongside the sparse and unregularized evidence maps, we also include visualizations from post-hoc methods.
 Figure 13: Class-specific explanation with the ResNet backbone.
@@
 Figure 14: Class-specific explanation with the VGG backbone.
-The application of our method to multi-class DR detection demonstrates the utility of class-specific explanations produced by the sparse SoftCAM, which more precisely highlight disease-relevant regions compared to the dense SoftCAM and the best-performing post-hoc method, ScoreCAM.
-In the example shown, the image is labeled as severe DR, and the highlighted regions correspond to suspicious areas, reflecting relevant DR lesions.
-F.5 Qualitative explanation on retinal OCT images
-For the multi-class tasks on retinal disease classification from OCT images, SoftCAM variants produced more focused and class-consistent explanations.
-In addition to the sparse and dense evidence maps, we also provide visualizations for GradCAM and Guided BP, as these were the best-performing post-hoc methods for the ResNet and VGG backbones, in terms of the Area Under the Deletion Curve (Tab.
-6).
+I.4 Qualitative explanation on retinal OCT images
+For multi-class OCT-based retinal disease classification, SoftCAM variants yielded more focused and class-consistent explanations.
+We also provide visualizations from other methods.
+Note that GradCAM and Guided BP were the best-performing post-hoc methods for ResNet and VGG according to the Area Under the Deletion Curve (Tab.
+7).
 Figure 15: Class-specific explanation with the ResNet backbone.
-SoftCAM applied to multi-class retinal disease classification demonstrates the utility of class-specific explanations, with the sparse SoftCAM variant more precisely highlighting disease-relevant regions compared to both the dense SoftCAM and the best-performing post-hoc methods, GradCAM and Guided Backpropagation.
-In the example shown, the image is labeled as Diabetic Macular Edema (DME), and the highlighted regions produced by sparse SoftCAM highlight suspicious areas, reflecting relevant retinal lesions.
+Sparse SoftCAM provides more precise class-specific localization than unregularized SoftCAM and leading post-hoc methods, highlighting clinically relevant regions.
 Figure 16: Class-specific explanation with the VGG backbone.
-SoftCAM applied to multi-class retinal disease classification demonstrates the utility of class-specific explanations, with the sparse SoftCAM variant more precisely highlighting disease-relevant regions compared to both the dense SoftCAM and the best-performing post-hoc methods, GradCAM and Guided Backpropagation.
-In the example shown, the image is labeled as Diabetic Macular Edema (DME), and the highlighted regions produced by sparse SoftCAM highlight suspicious areas, reflecting relevant retinal lesions.